### v3

# Visual Mamba: A Survey and New Outlooks

Abstract.
Mamba, a recent selective structured state space model, excels in long sequence modeling, which is vital in the large model era. Long sequence modeling poses significant challenges, including capturing long-range dependencies within the data and handling the computational demands caused by their extensive length. Mamba addresses these challenges by overcoming the local perception limitations of convolutional neural networks and the quadratic computational complexity of Transformers. Given its advantages over these mainstream foundation architectures, Mamba exhibits great potential to be a visual foundation architecture. Since January 2024, Mamba has been actively applied to diverse computer vision tasks, yielding numerous contributions. To help keep pace with the rapid advancements, this paper reviews visual Mamba approaches, analyzing nearly 300 papers. This paper begins by delineating the formulation of the original Mamba model. Subsequently, it delves into representative backbone networks, and applications categorized using different modalities, including image, video, point cloud, and multi-modal data. Particularly, we identify scanning techniques as critical for adapting Mamba to vision tasks, and decouple these scanning techniques to clarify their functionality and enhance their flexibility across various applications. Finally, we discuss the challenges and future directions, providing insights into new outlooks in this fast evolving area. A comprehensive list of visual Mamba models reviewed in this work is available at https://github.com/Ruixxxx/Awesome-Vision-Mamba-Models.
Mamba, State Space Model, Computer Vision, Application
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Artificial intelligence technologies, especially deep learning, have revolutionized numerous application fields. In the field of computer vision (CV), convolutional neural networks (CNNs) utilize local receptive fields and shared weights to process visual data, capitalizing on inductive biases such as locality and spatial invariance (Krizhevsky et al., 2012; Simonyan and Zisserman, 2015; He et al., 2016). Despite their efficient linear computational complexity with respect to image resolution and proficiency in modeling local patterns, CNNs have restricted receptive fields. This restriction limits their capability to capture larger spatial contexts, which is essential for comprehensively understanding scenes or complex spatial relations in tasks that demand global information. In recent years, Vision Transformers (ViTs) (Dosovitskiy et al., 2021), which utilize a self-attention mechanism (Vaswani et al., 2017) to process sequences of image patches, have demonstrated remarkable modeling capabilities across various visual tasks (Liu et al., 2021). Self-attention enables ViTs to capture long-range dependencies within images, providing a significant advantage over traditional CNNs that rely on local receptive fields. This capability allows ViTs to exhibit robust performance on varied datasets and scale effectively to large model sizes. However, the self-attention mechanism involves a quadratic computational cost to the number of patches, which limits the scalability of ViTs. The CV domain has long been dominated by CNNs and ViTs, each with their respective strengths and inherent limitations. To overcome their limitations, researchers have devoted significant effort to improving these models. Recently, structured state space models (Gu et al., 2022a, 2023) have garnered considerable attention due to their computational efficiency and principled capability in modeling long-range dependencies (Gu et al., 2020).
Figure 1. The statistics of Mamba-based papers released to date on vision tasks, spanning different modalities including Image, Video, Point Cloud, and Multi-Modal.
[…]
The state space model is a concept that is widely adopted in various disciplines. Its core idea is connecting the input and output sequences using a latent state. It takes different forms in different disciplines, such as Markov decision process in reinforcement learning (Hafner et al., 2020), dynamic causal modeling in computational neuroscience (Friston et al., 2003) and Kalman filters in controls (Kalman, 1960). Recently, the state space model (SSM) has been introduced to deep learning for sequence modeling and its parameters or mappings are learned by gradient descent (Gu et al., 2021). SSM is essentially a type of sequence transformation and can be incorporated into deep neural networks. It conceptually unifies the strengths of former paradigms for sequence model design, including continuous-time models (CTMs), recurrent neural networks (RNNs), and CNNs. Among these, although CNNs are primarily used for processing spatial data and are not sequential, they can be adapted for sequence modeling by leveraging their abilities to encode local contexts and facilitate parallelizable computations. However, SSMs have not been widely used in practice due to their extensive computational and memory requirements coming from the state representation. This situation has changed with the advent of the structured SSM (S4), which addresses these limitations through reparameterizing the state matrices (Gu et al., 2022a). Since then, a series of SSM transformations and neural network architectures incorporating SSM layers emerge (Gu et al., 2022a, b; Gupta et al., 2022; Orvieto et al., 2023; Smith et al., 2023; Gu et al., 2023; Hasani et al., 2023; Fu et al., 2023). However, SSMs’ constant sequence transitions restrict their context-based reasoning ability, which is important for the efficacy of models like Transformer (Vaswani et al., 2017). In Mamba-1 (Gu and Dao, 2024), the authors propose to address this by integrating a selection mechanism into the SSM, thus enabling the SSM to selectively propagate or forget information along the sequence or scan path based on the current token. Furthermore, to efficiently compute these selective SSMs, the authors develop a hardware-aware algorithm. In Mamba-2 (Dao and Gu, 2024a), the authors establish connections between the SSMs and variants of attention, and refine the selective SSM into a state space duality (SSD) algorithm to further improve computational efficiency. With the modeling power akin to Transformers and linear scalability in sequence length, Mamba becomes a promising foundation architecture for sequence modeling, which is crucial in the large model era.
Due to the growing adoption of techniques from sequence modeling or natural language processing (NLP) into CV, there is a rapid application of Mamba to CV tasks (Zhang et al., 2024i; Liu et al., 2024i; Heidari et al., 2024; Zou et al., 2024a). VMamba (Liu et al., 2024e), an early representative visual Mamba model, unfolds image patches into sequences along the horizontal and vertical dimensions of the image, and performs bi-directional scanning along these two directions. Another visual Mamba model Vim (Zhu et al., 2024c) leverages position embeddings to incorporate spatial information, inspired by ViT (Dosovitskiy et al., 2021). It also uses bi-directional SSM for handling the non-causal image sequences. Similarly, several other notable studies (Li et al., 2024h; Yang et al., 2024a; Patro and Agneeswaran, 2024b; Huang et al., 2024d; Pei et al., 2024) delve into the exploration of visual backbone networks, consistently achieving competitive performance across classification, detection, and segmentation tasks. To highlight the balance between efficiency and effectiveness in visual Mamba models, Fig. 5 offers a graphical representation contrasting their performance with computational complexity. In addition to these efforts, Mamba has been applied across a diverse range of vision modalities and their respective applications, encompassing image processing, video analysis, point cloud processing, and multi-modal scenarios. Endowed with the aforementioned modeling capability and linear scalability, Mamba stands out as a promising foundation architecture for CV tasks. The growing interest among researchers in applying Mamba to various vision tasks is reflected in the increasing number of studies dedicated to this exploration, as plotted in Fig. 1.
In the rapidly evolving field of CV, Mamba has emerged as a significant advancement. Keeping pace with the latest research is critical for the community. Therefore, this paper aims to provide a comprehensive review of the applications of Mamba in vision tasks, shedding light on both its foundational elements and diverse applications across various modalities. Our contributions are summarized as follows:
  1. 1.
    Formulation of Mamba (Section 2): We provide an introductory overview of the operational principles of the Mamba (Gu and Dao, 2024; Dao and Gu, 2024a) and highlight its distinctions from traditional state space models.
  2. 2.
    Backbone Networks (Section 3): We provide a detailed examination of several representative visual Mamba backbone networks. This analysis aims to elucidate the core principles and innovations that underpin the visual Mamba framework.
  3. 3.
    Applications (Section 4): We categorize applications of Mamba by different modalities, such as image, video, point cloud, and multi-modal data. Each category is explored in depth to highlight how the Mamba architecture adapts to and benefits individual modalities.
  4. 4.
    Challenges and Future Directions (Section 5): We examine the challenges of visual Mamba models, focusing on their scalability, causality, in-context learning, and trustworthiness. More importantly, we outline prospective directions for visual Mamba models, providing new outlooks into exploring their untapped potential.
2. Formulation of Mamba
Mamba (Gu and Dao, 2024) is a recent sequence model aiming at improving the context-based reasoning ability of state space model (SSM) by simply making its parameters to be functions of the input. The SSM here especially refers to the sequence transformation used in the structured state space sequence model (S4) (Gu et al., 2022a), which can be incorporated into deep neural networks. Mamba-2 (Dao and Gu, 2024b) further provides a theoretical framework that connects structured SSMs and variants of attention, thereby enabling the transfer of algorithmic and system optimizations originally developed for Transformers to SSMs. In the following, we elaborate on the core concepts of Mamba-1 and Mamba-2.
2.1. State Space Model
The SSM transformation in S4 (Gu et al., 2022a) originates from the classical state space model, which maps a 1D input signal $x(t)\in\mathbb{R}$ to a 1D output signal $y(t)\in\mathbb{R}$ through an N-D latent state $h(t)\in\mathbb{R}^{N}$. This transformation is mathematically formulated as linear ordinary differential equations (ODEs):
(1) $h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t),$
where $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the evolution parameter and $\mathbf{B}\in\mathbb{R}^{N\times 1},\mathbf{C}\in\mathbb{R}^{1\times N}$ are the projection parameters of neural networks in deep learning. The term $h^{\prime}(t)$ denotes the derivative of $h(t)$ with respect to time $t$.
To deal with the discrete input sequence $\boldsymbol{x}=(x_{0},x_{1},...)\in\mathbb{R}^{L}$, various discretization rules can be used to discretize the parameters in Eq. (1) using a step size $\Delta$, which can be seen as the resolution of the continuous input $x(t)$. Following previous work (Tustin, 1947), S4 (Gu et al., 2022a) discretizes these parameters using the bilinear method. Following (Gupta et al., 2022), Mamba (Gu and Dao, 2024) employs the zero-order hold (ZOH) assumption [^1] ZOH assumes that the sample value of $x$ remains constant over each sampling interval $\Delta$. For more details on ZOH, refer to https://en.wikipedia.org/wiki/Zero-order_hold to solve the ODEs. Specifically, this results in the conversion of the continuous parameters $\mathbf{A},\mathbf{B}$ into their discrete counterparts $\overline{\mathbf{A}},\overline{\mathbf{B}}$ as follows:
(2) $\overline{\mathbf{A}}=\exp(\Delta\mathbf{A}),\quad\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}.$
After discretizing $\mathbf{A},\mathbf{B}$ to $\overline{\mathbf{A}},\overline{\mathbf{B}}$, the Eq. (1) can be reformulated as:
(3) $h_{t}=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t},\quad y_{t}=\mathbf{C}h_{t}.$
Eq. (3) represents a sequence-to-sequence mapping from $x_{t}$ to $y_{t}$. This configuration allows the discretized SSM to be computed as RNN. However, due to its sequential nature, this discretized recurrent SSM is impractical for training.
For efficient parallelizable training, this recursive process can be reformulated and computed as a convolution (Gu et al., 2022a):
(4) $\overline{\mathbf{K}}=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{AB}},\ldots,\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}),\quad\boldsymbol{y}=\boldsymbol{x}*\overline{\mathbf{K}},$
where $L$ denotes the length of the input sequence $\boldsymbol{x}$ and * represents the convolution operation. The vector $\overline{\mathbf{K}}\in\mathbb{R}^{L}$ is the SSM convolution kernel, enabling the simultaneous synthesis of outputs across the sequence. Given $\overline{\mathbf{K}}$, the convolution operation in Eq. (4) can be efficiently computed using the fast Fourier transforms (FFTs).
2.2. Mamba-1: Selective SSM
As seen, the parameters in SSM indicated by either Eq. (1), Eq. (3) or Eq. (4) remain invariant with respect to the input or temporal dynamics. Mamba (Gu and Dao, 2024) identifies this linear time-invariant (LTI) property as a fundamental limitation of SSM when it comes to context-based reasoning. To address this issue, Mamba incorporates a selection mechanism. The selection mechanism is implemented by simply configuring the parameters of SSM as functions of the input, thus achieving input-dependent interactions along the sequence. Specifically, parameters $\mathbf{B},\mathbf{C},\Delta$ are dependent on the input sequence $\boldsymbol{x}$:
(5) $\mathbf{B}=s_{B}(\boldsymbol{x}),\quad\mathbf{C}=s_{C}(\boldsymbol{x}),\quad\Delta=\tau_{\Delta}(Parameter+s_{\Delta}(\boldsymbol{x})),$
where $s_{B}(\boldsymbol{x})=Linear_{N}(\boldsymbol{x})$ and $s_{C}(\boldsymbol{x})=Linear_{N}(\boldsymbol{x})$ both project the input into a dimension $N$. $s_{\Delta}(\boldsymbol{x})=Broadcast_{D}(Linear_{1}(\boldsymbol{x}))$ first projects the input to dimension 1, and then broadcasts it to dimension $D$. $\tau_{\Delta}$ is the $softplus$ function. Here we present the complete shape of $\boldsymbol{x}\in\mathbb{R}^{B\times L\times D}$, where $B$ denotes the batch size, $L$ is the sequence length, and $D$ is the number of channels. Accordingly, the parameters $\mathbf{B}$ and $\mathbf{C}$ are each shaped as $\mathbb{R}^{B\times L\times N}$, and $\Delta$ is shaped as $\mathbb{R}^{B\times L\times D}$.
The resulting selective SSM cannot be computed as either RNN or convolution. Mamba (Gu and Dao, 2024) employs a hardware-aware algorithm to efficiently compute the selective SSM. The hardware-aware algorithm leverages three classical techniques: parallel scan, kernel fusion, and recomputation.
2.3. Mamba-2: State Space Duality
To leverage the collective advancements made by the community on Transformers, Mamba-2 (Dao and Gu, 2024a) first reformulates the SSM transformation through induction as:
(6) $y_{t}=\sum_{s=0}^{t}\mathbf{C}_{t}^{\top}\mathbf{A}_{t:s}^{\times}\mathbf{B}_{s}x_{s},$
where $\mathbf{A}_{t:s}^{\times}=\mathbf{A}_{t}\times\cdots\times\mathbf{A}_{s+1},s<t$. The matrix transformation form of SSM can then be derived as:
(7) $\boldsymbol{y}=\text{SSM}(\mathbf{A},\mathbf{B},\mathbf{C})(\boldsymbol{x})=\mathbf{M}\boldsymbol{x},$
where $\mathbf{M}_{ji}=\mathbf{C}_{j}^{\top}\mathbf{A}_{j}\cdots\mathbf{A}_{i+1}\mathbf{B}_{i}$. Consequently, SSM also has a quadratic algorithm besides the linear algorithm.
Mamba-2 considers a special case of structured SSM when $\mathbf{A}_{j}$ is a scalar. Then the matrix can be rearranged as:
(8) $\mathbf{M}_{ji}=\mathbf{A}_{j:i}\cdot(\mathbf{C}_{j}^{\top}\mathbf{B}_{i}).$
This can be vectorized into:
(9) $\mathbf{M}=\mathbf{L}\circ(\mathbf{C}\mathbf{B}^{\top}),$
which is consistent with the original definition of masked kernel attention. To sum up, SSM with a scalar-identity structure in the $\mathbf{A}$ matrix, and structured masked attention with a 1-semiseparable structure in the $\mathbf{L}$ mask are dual to each other, exhibiting identical linear and quadratic forms.
Mamba-2 introduces an SSD algorithm based on block decompositions of $\mathbf{M}$, leveraging both the linear SSM recurrence and its quadratic dual form to improve computational and memory efficiency while utilizing matrix multiplication units on modern hardware. This SSD algorithm achieves a $2$-8×8$\times$8 × speedup over the hardware-aware Mamba-1 implementation.
2.4. Mamba Architecture
Mamba-1 is a simplified SSM architecture, drawing insights from the H3 architecture (Fu et al., 2023), which serves as the basis of many SSM architectures. Specifically, the H3 (Fu et al., 2023) block integrates two discrete SSMs to explicitly simulate linear attention. The commonly used SSM architectures alternately stack the H3 block (Fu et al., 2023) and the MLP block prevalent in modern neural networks. Unlike them, Mamba-1 integrates these two blocks to construct the simplified Mamba-1 block, as illustrated in Fig. 2. The Mamba-1 block can be viewed from two distinct perspectives. Firstly, it replaces the first multiplicative gate in the linear attention-like or H3 (Fu et al., 2023) block with an activation function. Secondly, it incorporates the SSM transformation into the primary pathway of the MLP block. The overall architecture of Mamba-1 consists of repeated Mamba-1 blocks interleaved with standard normalization layers and residual connections.
Figure 2. Mamba-1 block: sequential generation of SSM parameters vs. Mamba-2 block: parallel generation of SSM parameters.
[…]
Different from Mamba-1, which maps from $\boldsymbol{x}$ to $\boldsymbol{y}$ under the SSM point of view and sequentially computes $\mathbf{A},\mathbf{B},\mathbf{C}$ after $\boldsymbol{x}$, Mamba-2 produces $\mathbf{A},\boldsymbol{x},\mathbf{B},\mathbf{C}$ in parallel as the SSD layer is regarded as a mapping from $\mathbf{A},\boldsymbol{x},\mathbf{B},\mathbf{C}$ to $\boldsymbol{y}$. Thus the Mamba-2 block simplifies the Mamba-1 block by removing sequential linear projections; namely, SSM parameters $\mathbf{A},\mathbf{B},\mathbf{C}$ are generated at the beginning of the block, independent of the SSM input $\boldsymbol{x}$, as illustrated in Fig. 2. Additionally, a normalization layer as in NormFormer (Shleifer et al., 2021) is added after the multiplicative gate branch to enhance stability.
Mamba inherits the linear scalability in sequence length of the state space models, and also realizes the modeling ability of Transformers. Mamba exhibits the significant advantages of two primary types of foundation architectures in CV, i.e. CNNs and Transformers, making it a promising foundation architecture for CV. In contrast to Transformers, which rely on explicitly storing the entire context for context-based reasoning, Mamba utilizes a selection mechanism, which selectively propagates or filters information along the sequence based on the current token. This ensures the modeling process is influenced solely by the past and current inputs along the sequence, thereby adhering to the principles of causality. Therefore, adapting the inherent 1D and causal characteristics of this selection mechanism to align with the nature of visual data has become a primary focus for researchers applying Mamba to CV.
3. Visual Mamba Backbone Networks
This section reviews representative visual Mamba backbone networks, categorizing them by key architectural advancements and optimization strategies. A comparative analysis of their performance is also presented.
Figure 3. Scanning techniques, categorized into four groups, i.e., scan direction, scan axis, scan continuity, and scan sampling.
[…]
3.1. Architectural Advancements
The typical architecture of sequence models first transforms 2D images into sequences and then applies the sequence transformations. Following this, the architectural advancements of visual Mamba backbone networks focus on tokenization methods, scanning techniques, and architecture designs.
3.1.1. Tokenization
The 2D images are first converted into sequences of visual tokens via a stem module, typically comprising a convolution layer followed by a linear projection layer, to be processed. Given the inherent causal properties of SSM operations, adding positional embeddings to visual tokens is optional.
Some approaches treat these visual tokens as 1D structures, applying 1D SSM transformations and 1D convolution operations, as demonstrated in Vim (Zhu et al., 2024c). Other approaches preserve the 2D structure of visual tokens and adapt operations in Mamba-based blocks to 2D, as seen in VMamba (Liu et al., 2024e). In addition to visual tokens, other token types can be concatenated or added. Some works insert extra tokens to act as registers (Darcet et al., 2024) or to perceive discontinuities between rows or columns.
Differently, GlobalMamba (Wang et al., 2024p) segments images into multiple frequency bands and spatially downsamples them based on their frequency ranges, with lower resolutions for bands in smaller frequency ranges. These bands are then tokenized into causal sequences using a lightweight CNN. Through this, the 2D images are transformed into sequences of causal tokens containing global information.
3.1.2. Scanning Techniques
The selective scanning mechanism is the key component of Mamba. However, its original design for 1D causal sequences poses challenges when adapting it to non-causal visual data. Significant research efforts have focused on developing scanning techniques to overcome these challenges. To provide a clear understanding, we categorize these scanning techniques into four main groups: scan direction, scan axis, scan continuity, and scan sampling, forming a decoupled taxonomy. This categorization also reveals different objectives of the scanning techniques. The scan direction addresses the non-causal characteristics of visual sequences; the scan axis deals with the high dimensionality inherent in visual data; the scan continuity considers the spatial continuity of the patches along the scanning path; the scan sampling divides the full image into sub-images to capture spatial information. The illustration of these four groups is shown in Fig. 3.
Scan Direction
After unfolding the visual data into sequences, different scan directions can be employed to handle the non-causal nature of these sequences.
Single-directional scanning (SD) represents the original selective scanning technique in Mamba (Gu and Dao, 2024), which processes sequences in a single direction along sequence length. Bi-directional scanning (BD) processes the sequences from the forward and backward directions, thus enhancing the receptive fields reciprocally.
Scan Axis
Visual data differ from typical sequences by possessing 2D or higher-dimensional spatial information that encapsulates both local and global contexts. Current approaches unfold visual tokens along various axes of the original visual data to thoroughly integrate this spatial information. For instance, the scanning axes for a 2D image typically include the horizontal (H), vertical (V), left diagonal (LD), and right diagonal (RD). For 3D visual data, the axes also involve the depth (D) dimension. Additionally, Vim-F (Zhang et al., 2024a) applies the Fourier Transform to convert the image into the frequency domain (F), enabling a global receptive field during the scanning.
Beyond the spatial axes, certain approaches apply scanning along the channel (C) dimension to capture additional channel-specific information.
Scan Continuity
Alternative techniques are employed to handle the 2D spatial information in visual data.
Raster scanning flattens images into sequences following the layout of computer memory. However, this technique can cause spatial discontinuities between adjacent rows or columns, as spatial neighboring tokens in the image may become separated in the sequence. In contrast, Zigzag scanning serializes images following a zigzag path, alternating between left-to-right and right-to-left directions across rows (or columns). This ensures continuous scanning by maintaining spatial continuity among adjacent tokens. Some approaches utilize fractal scanning curves to better capture the structural information within the image. Specifically, Hilbert scanning leverages the Hilbert curve, a space-filling curve that recursively subdivides the image and traverses its regions in a structured manner. This technique effectively preserves the spatial relationships between tokens, significantly enhancing their adjacency within the sequence. Reordering scanning divides the sequence into non-overlapping segments and successively samples instances from each segment in a fixed order, ensuring systematic coverage of the entire sequence. Differently, Shuffle scanning randomizes the token sequence order to enhance positional transformation invariance.
Scan Sampling
In addition to operating on the full image, referred to as Global Sampling, various other sampling techniques are employed for specific purposes.
Local Sampling divides the image into sub-images and performs scanning within each, improving the model’s ability to capture local dependencies. Atrous Sampling samples the image into sub-images with skipping, capturing global dependencies while reducing computational complexity. Multi-scale Sampling uses depthwise convolution to generate multi-scale feature maps for scanning. The shorter sequences of low-resolution feature maps help mitigate the long-range forgetting problem while reducing computational cost. Quadtree Sampling uses a quadtree-based image partitioning strategy that estimates the locality score of each token based on its features. Tokens are then adaptively divided into window quadrants at both coarse and fine granularities, enabling the capture of local dependencies across multiple scales. Channel Sampling divides the image channels into groups to improve computational efficiency. Notably, within these sub-regions, different combinations of the former three groups of scanning techniques can be applied, allowing for diverse processing approaches without the necessity for uniformity across all sub-regions.
These four groups of scanning techniques are interoperable and can be synergistically combined to enhance visual data analysis.
3.1.3. Architecture
The core components of the visual Mamba backbone network architecture are various blocks, formed by different combinations of the previously discussed tokenization methods, scanning techniques, and the selective SSM transformation.
These network architectures can consist of pure Mamba blocks that solely rely on the Mamba architecture for processing visual data. Vim (Zhu et al., 2024c) is a pure Mamba network structured as a series of identical Vim blocks, which is a Mamba block integrating a backward SSM path alongside the forward one. VMamba (Liu et al., 2024e) is another pure Mamba network. Its block is similar to the Mamba block but replaces the 1D convolution layer with a 2D depth-wise convolution layer and the selective SSM with 2D-Selective-Scan (SS2D). SS2D unfolds the input patches into sequences along four distinct scanning paths (Cross-Scan), processes each sequence using separate selective SSM transformations in parallel, and then reshapes and combines the resulting sequences to obtain the final output map (Cross-Merge). The PlainMamba (Yang et al., 2024a) block is similar to the VMamba block, with two key differences. Firstly, it employs the continuous zigzag scanning technique to maintain spatial adjacency among tokens and prevent discontinuities. Secondly, a direction-aware updating technique is introduced to explicitly incorporate relative 2D position information into the selective scanning process. Differently, Mamba-ND (Li et al., 2024h) treats the 1D Mamba layer as a black box and explores how to unravel and order multi-dimensional data. Its results show that a chain of Mamba layers and simple alternating-directional orderings achieve superior performance.
Figure 4. Visual Mamba blocks, including Vision Mamba (Vim) (Zhu et al., 2024c), Visual State Space (VSS) (Liu et al., 2024e), MSVMamba (Shi et al., 2024a), PlainMamba (Yang et al., 2024a), and LocalMamba (Huang et al., 2024d) blocks. The original Mamba (Gu and Dao, 2024) block is presented as a reference for the advancements in the visual Mamba blocks. The scanning techniques and their decoupling results are displayed to the left and above the corresponding blocks, respectively.
[…]
Table 1. Architectural Advancements for Visual Mamba Backbone Networks.
Methods Tokenization Scan Strategy Architecture Venue Code
Direction Axis Continuity Sampling Block Structure
Vim (Zhu et al., 2024c) 1D BD H Raster Global Pure Plain ICML24
VMamba (Liu et al., 2024e) 2D BD H/V Raster Global Pure Hierarchical NeurIPS24
Mamba-ND (Li et al., 2024h) 1D BD H/V/D Raster Global Pure Plain ECCV24 Oral
LocalMamba (Huang et al., 2024d) 1D BD H/V Raster Local Hybrid Plain/Hierarchical
EfficientVMamba (Huang et al., 2024d) 2D BD H/V Raster Atrous Hybrid Hierarchical
SiMBA (Patro and Agneeswaran, 2024b) 1D SD H Raster Global Hybrid Plain
PlainMamba (Yang et al., 2024a) 2D BD H/V Zigzag Global Pure Plain BMVC24
MSVMamba (Shi et al., 2024a) 2D BD H/V Raster Multi-scale Hybrid Hierarchical NeurIPS24
FractalMamba (Tang et al., 2024c) 2D BD H/V Hilbert Global Pure Plain
Mamba® (Wang et al., 2024j) 1D + Register BD H Raster Global Pure Plain
Vim-F (Zhang et al., 2024a) 1D BD H + F Raster Global Pure Plain
MambaVision (Hatamizadeh and Kautz, 2024) 1D SD H Raster Global Hybrid Hierarchical
GroupMamba (Shaker et al., 2024) 1D BD H/V Zigzag Channel Hybrid Hierarchical
ShuffleMamba (Huang et al., 2024b) 1D BD H Raster/Shuffle Global Pure Plain
StableMamba (Suleman et al., 2024) 1D BD H Raster Global Hybrid Plain
QuadMamba (Xie et al., 2024c) 2D BD H/V Raster Quadtree Pure Hierarchical NeurIPS24
GlobalMamba (Wang et al., 2024p) F-based Global BD H/V Raster Global Pure Hierarchical
The network architectures can also incorporate hybrid Mamba blocks that combine the Mamba architecture with other neural network architectures, such as CNNs and attentions, to capitalize on their complementary advantages. MSVMamba (Shi et al., 2024a) incorporates a convolutional feed-forward network in each block to enhance channel-wise information exchange and local feature extraction. LocalMamba (Huang et al., 2024d) introduces a spatial and channel attention module before patch merging to enhance the integration of directional features and reduce redundancy. SiMBA (Patro and Agneeswaran, 2024b) uses Mamba for sequence modeling and introduces a new channel modeling technique based on Fourier Transforms.
Fig. 4 illustrates a suite of pure and hybrid visual Mamba blocks, including Vision Mamba (Vim) (Zhu et al., 2024c), Visual State Space (VSS) (Liu et al., 2024e), MSVMamba (Shi et al., 2024a), PlainMamba (Yang et al., 2024a), and LocalMamba (Huang et al., 2024d) blocks. The Mamba (Gu and Dao, 2024) block is also included to facilitate a direct comparison, highlighting the evolutionary design of these blocks in the visual domain. The detailed illustrations of scanning techniques of different blocks and their decoupled results presented in Fig. 4 also validate the logic behind our categorization of scanning techniques.
Table 2. Optimization Strategies for Visual Mamba Backbone Networks.
Methods Strategy Venue Code
Mamba-prune (Zhan et al., 2024) Token Pruning NeurIPS24
VSSD (Shi et al., 2024b) Visual State Space Duality
V2M (Wang et al., 2024m) 2D SSM
Spatial-Mamba (Xiao et al., 2024b) Structure-aware State Fusion
SparX-Mamba (Lou et al., 2024) Cross-layer Connection
HRVMamba (Zhang et al., 2024g) Multi-resolution Parallel Design
ARM (Ren et al., 2024a) Autoregressive Pre-training
MAP (Liu and Yi, 2024) Masked Autoregressive Pre-training
Built on different types of blocks, existing architectures also differ in their use of plain or hierarchical structures. Plain structures typically stack identical blocks, while hierarchical structures introduce down-sampling layers between block stacks to generate hierarchical representations.
Details of tokenization methods, scanning techniques, and architecture designs of visual Mamba backbone networks are outlined in Table. 1.
3.2. Optimization Strategies
In addition to architectural advancements, researchers also explore additional optimization strategies to enhance the efficiency and performance of visual Mamba backbone networks. A summary of these strategies is presented in Table. 2. Mamba-prune (Zhan et al., 2024) introduces token pruning into Mamba-based networks for accelerating the computation. VSSD (Shi et al., 2024b) adapts SSD to a non-causal mode to better handle the non-causality of visual data, while V2M (Wang et al., 2024m) and Spatial-Mamba (Xiao et al., 2024b) extend the SSM to 2D forms. SparX-Mamba (Lou et al., 2024) and HRVMamba (Zhang et al., 2024g) introduce additional designs for cross-layer feature aggregation and dense prediction, respectively. ARM (Ren et al., 2024a) and MAP (Liu and Yi, 2024) focus on improving the model performance via pre-training.
Table 3. Comparison of different backbones on ImageNet-1K (Deng et al., 2009) classification at the resolution of $224^{2}$ (Part I). ‘Conv’, ‘Attn’, and ‘Mamba’ denote architectures based on CNN, Transformer, and Mamba, respectively. indicates that the backbone has undergone additional training using the Token Labeling objective (Wang et al., 2022a). indicates results are obtained with MESA (Du et al., 2022). indicates results reproduced by DeiT (Touvron et al., 2021). indicates hierarchical structures while others are plain structures. l represents local sampling. The best results are marked with bold and underline, the second-best with bold only, and the third-best with underline only.
[…]
Figure 5. Comparative analysis of the performance and computational complexity across various visual backbone architectures, encompassing Convolution-based methods (Radosavovic et al., 2020; Liu et al., 2022; Yu and Wang, 2024), Transformer-based methods (Touvron et al., 2021; Liu et al., 2021; Yao et al., 2022; Patro and Agneeswaran, 2023; Yuan et al., 2023), and Mamba-based methods (Zhu et al., 2024c; Yang et al., 2024a; Liu et al., 2024e; Shi et al., 2024a; Wang et al., 2024j; Xie et al., 2024c; Shi et al., 2024b). The symbol size is proportional to the parameter count of respective model, providing a visual indicator of model scale and complexity. Note that all data are accessed from released academic papers to ensure fairness and credibility.
[…]
3.3. Experimental Results
In this section, we compare the experimental results of different visual Mamba backbone networks with established architectures based on CNNs or Transformers on standard computer vision benchmarks.
3.3.1. Image Classification
Image classification performance is compared on ImageNet-1K (Deng et al., 2009) and the Top-1 accuracy results from the respective papers are listed in Table. 3. To clearly compare different networks, we also plot their performance and computational complexity in Fig. 5. As seen, the performance of all the models improves as model sizes increase, illustrating the trend of enhancing performance through model scaling. It is evident that with similar computational complexity, all visual Mamba networks outperform the CNN-based RegNetY (Radosavovic et al., 2020) and the Transformer-based DeiT (Touvron et al., 2021). Furthermore, many visual Mamba networks surpass the CNN-based MambaOut (Yu and Wang, 2024) and Transformer-based Swin Transformer (Liu et al., 2021). These results demonstrate the superior performance of visual Mamba networks compared to advanced CNNs and common Transformers. However, Transformer-based networks are highly competitive, particularly SVT (Patro and Agneeswaran, 2023), Wave-ViT (Yao et al., 2022), and VOLO (Yuan et al., 2023), which consistently demonstrate superior performance over other networks. Notably, VOLO (Yuan et al., 2023) also exhibits excellent scalability. Currently, despite their efficiency, most Mamba-based networks are limited to small-scale implementations with FLOPs under 21G. While Mamba® (Wang et al., 2024j) and ARM (Ren et al., 2024a) have demonstrated scalability to larger models, there remains a need for performance enhancements. To sum up, visual Mamba networks exhibit promising performance but still underperform advanced Transformer-based networks. Further exploration is also needed to scale visual Mamba networks to larger configurations.
3.3.2. Object Detection
Object detection and instance segmentation performance are compared on the MS COCO (Lin et al., 2014) via Mask R-CNN (He et al., 2017) and the results from the respective papers are listed in Table. 4. The object detection and instance segmentation performance of all the models follows the scaling law as well. It can be observed that with comparable computational complexity, visual Mamba networks surpass the CNN-based ConvNeXt (Liu et al., 2022), except for EfficientVMamba (Pei et al., 2024), which is designed for lightweight purposes. Most visual Mamba networks demonstrate superior performance compared to all CNN-based architectures and a majority of Transformer-based models. Nonetheless, in certain scenarios, their performance is slightly below that of SG-Former (Ren et al., 2023). These results demonstrate that Mamba’s ability to capture long-range dependencies and utilize dynamic weights is advantageous for dense prediction tasks. Currently, Mamba-based models are not specifically tailored for these tasks, indicating potential for future research to explore and enhance their performance.
3.3.3. Semantic Segmentation
Semantic segmentation performance is compared on ADE20K (Zhou et al., 2019) utilizing UperNet (Xiao et al., 2018) and the results from the respective papers are listed in Table. 5. The semantic segmentation performance of different networks is similar to their object detection performance. Many visual Mamba networks outperform all CNN-based networks and most Transformer-based networks. However, their performance lags behind that of SG-Former (Ren et al., 2023). These results further demonstrate Mamba’s efficacy in dense prediction tasks. Further exploration of Mamba-based models specifically designed for these tasks is warranted.
Table 4. Results of object detection and instance segmentation on MS COCO (Lin et al., 2014) $mini$-$val$ using Mask R-CNN (He et al., 2017) 1$\times$ schedule. ‘Conv’, ‘Attn’, and ‘Mamba’ denote architectures based on CNN, Transformer, and Mamba, respectively. FLOPs are computed using an input size 1280$\times$800. indicates FLOPs computed using an input size 1333$\times$800. indicates hierarchical structures while others are plain structures. The best results are marked with bold and underline, the second-best with bold only, and the third-best with underline only.
Backbone Type Params FLOPs $\mathbf{AP^{b}}$ $\mathbf{AP^{b}_{50}}$ $\mathbf{AP^{b}_{75}}$ $\mathbf{AP^{m}}$ $\mathbf{AP^{m}_{50}}$ $\mathbf{AP^{m}_{75}}$
ConvNeXt-T (Liu et al., 2022) Conv 48.0 M 262.0 G 44.2 66.6 48.3 40.1 63.3 42.8
MambaOut-T (Yu and Wang, 2024) Conv 43.0 M 262.0 G 45.1 67.3 49.6 41.0 64.1 44.1
Swin-T (Liu et al., 2021) Attn 48.0 M 264.0 G 42.2 - - 39.1 - -
PVTv2-B2 (Wang et al., 2022b) Attn 45.0 M - 45.3 67.1 49.6 41.2 64.2 44.4
ViL-S (Zhang et al., 2021) Attn 45.0 M 218.3 G 44.9 67.1 49.3 41.0 64.2 44.1
SG-Former-S (Ren et al., 2023) Attn 41.0 M - 47.4 69.0 52.0 42.6 65.9 46.0
[…]
VMamba-T (Liu et al., 2024e) Mamba 50.0 M 270.0 G 47.4 69.5 52.0 42.7 66.3 46.0
FractalMamba-T (Tang et al., 2024c) Mamba 50.0 M 270.0 G 47.8 70.0 52.4 42.9 66.6 46.3
EffVMamba-S (Pei et al., 2024) Mamba 31.0 M 197.0 G 39.3 61.8 42.6 36.7 58.9 39.2
V2M-S (Wang et al., 2024m) Mamba 50 M - 47.6 69.4 52.2 42.9 66.5 46.3
GlobalMamba-T (Wang et al., 2024p) Mamba 50 M - 47.6 69.4 52.5 42.9 66.5 46.0
QuadMamba-T (Xie et al., 2024c) Mamba 30 M 213 G 42.3 64.6 46.2 38.8 61.6 41.4
MSVMamba-M (Shi et al., 2024a) Mamba 32.0 M 201.0 G 43.8 65.8 47.7 39.9 62.9 42.9
LocalVMamba-T (Huang et al., 2024d) Mamba 45.0 M 291.0 G 46.7 68.7 50.8 42.2 65.7 45.5
GroupMamba (Shaker et al., 2024) Mamba 40 M 279 G 47.6 69.8 52.1 42.9 66.5 46.3
Spatial-Mamba-T (Xiao et al., 2024b) Mamba 46 M 261 G 47.6 69.6 52.3 42.9 66.5 46.2
VSSD-M (Shi et al., 2024b) Mamba 33 M 220 G 45.4 67.5 49.8 41.3 64.5 44.6
VSSD-T (Shi et al., 2024b) Mamba 44 M 265 G 46.9 69.4 51.4 42.6 66.4 45.9
ConvNeXt-S (Liu et al., 2022) Conv 70.0 M 348.0 G 45.4 67.9 50.0 41.8 65.2 45.1
MambaOut-S (Yu and Wang, 2024) Conv 65.0 M 354.0 G 47.4 69.1 52.4 42.7 66.1 46.2
ViT-Adpt-S (Chen et al., 2023) Attn 47.8 M - 44.7 65.8 48.3 39.9 62.5 42.8
Swin-S (Liu et al., 2021) Attn 69.0 M 354.0 G 44.8 - - 40.9 - -
PVTv2-B3 (Wang et al., 2022b) Attn 64.9 M - 47.0 68.1 51.7 42.5 65.7 45.7
ViL-M (Zhang et al., 2021) Attn 60.1 M 293.8 G 47.6 69.8 52.1 43.0 66.9 46.6
SG-Former-M (Ren et al., 2023) Attn 51.0 M - 48.2 70.3 53.1 43.6 66.9 47.0
[…]
VMamba-S (Liu et al., 2024e) Mamba 64.0 M 357.0 G 48.7 70.0 53.4 43.7 67.3 47.0
EffVMamba-B (Pei et al., 2024) Mamba 53.0 M 252.0 G 43.7 66.2 47.9 40.2 63.3 42.9
V2M-B (Wang et al., 2024m) Mamba 70 M - 48.9 70.2 53.6 43.8 67.5 47.1
GlobalMamba-S (Wang et al., 2024p) Mamba 70 M - 49.0 70.5 53.5 43.9 67.5 47.0
QuadMamba-S (Xie et al., 2024c) Mamba 55 M 301 G 46.7 69.0 51.3 42.4 65.9 45.6
MSVMamba-T (Shi et al., 2024a) Mamba 53.0 M 252.0 G 46.9 68.8 51.4 42.2 65.6 45.4
LocalVMamba-S (Huang et al., 2024d) Mamba 69.0 M 414.0 G 48.4 69.9 52.7 43.2 66.7 46.5
SiMBA-S (Patro and Agneeswaran, 2024b) Mamba 60.0 M 382.0 G 46.9 68.6 51.7 42.6 65.9 45.8
Spatial-Mamba-S (Xiao et al., 2024b) Mamba 63 M 315 G 49.2 70.8 54.2 44.0 67.9 47.5
VSSD-S (Shi et al., 2024b) Mamba 59 M 325 G 48.4 70.1 53.1 43.5 67.2 47.1
MambaOut-B (Yu and Wang, 2024) Conv 100.0 M 495.0 G 47.4 69.3 52.2 43.0 66.4 46.3
ViT-Adpt-B (Chen et al., 2023) Attn 120.2 M - 47.0 68.2 51.4 41.8 65.1 44.9
PVTv2-B4 (Wang et al., 2022b) Attn 82.2 M - 47.5 68.7 52.0 42.7 66.1 46.1
PVTv2-B5 (Wang et al., 2022b) Attn 101.6 M - 47.4 68.6 51.9 42.5 65.7 46.0
Swin-B (Liu et al., 2021) Attn 107.0 M 496.0 G 46.9 - - 42.3 - -
ViL-B (Zhang et al., 2021) Attn 76.1 M 384.4 G 48.6 70.5 53.4 43.6 67.6 47.1
SG-Former-B (Ren et al., 2023) Attn 95.0 M - 49.2 70.6 54.3 68.2 68.1 47.7
[…]
VMamba-B (Liu et al., 2024e) Mamba 108.0 M 485.0 G 49.2 70.9 53.9 43.9 67.7 47.6
GlobalMamba-B (Wang et al., 2024p) Mamba 108 M - 49.3 71.4 54.2 44.2 68.4 47.7
Spatial-Mamba-B (Xiao et al., 2024b) Mamba 115 M 494 G 50.4 71.8 55.3 45.1 69.1 49.1
Table 5. Results of semantic segmentation on ADE20K (Zhou et al., 2019) $val$ at the resolution of $512^{2}$ using UperNet (Xiao et al., 2018). ‘Conv’, ‘Attn’, and ‘Mamba’ denote architectures based on CNN, Transformer, and Mamba, respectively. FLOPs are calculated using an input size 512$\times$2048. ‘SS’ and ‘MS’ denote single-scale and multi-scale testing, respectively. MLN: multi-level neck. indicates hierarchical structures while others are plain structures. The best results are marked with bold and underline, the second-best with bold only, and the third-best with underline only.
Backbone Type Params FLOPs mIoU (SS) mIoU (MS)
ViT-Adpt-T (Chen et al., 2023) Attn 36.1 M - 42.6 43.6
Vim-Ti (Zhu et al., 2024c) Mamba 13.0 M - 41.0 -
PlainMamba-L1 (Yang et al., 2024a) Mamba 35.0 M 174.0 G 44.1 -
EffVMamba-T (Pei et al., 2024) Mamba 14.0 M 230.0 G 38.9 39.3
QuadMamba-T (Xie et al., 2024c) Mamba 40 M 886 G 44.3 45.1
VSSD-M (Shi et al., 2024b) Mamba 42 M 893 G 45.6 46.0
ConvNeXt-T (Liu et al., 2022) Conv 60.0 M 939.0 G 46.0 46.7
MambaOut-T (Yu and Wang, 2024) Conv 54.0 M 938.0 G 47.4 48.6
ViT-Adpt-S (Chen et al., 2023) Attn 57.6 M - 46.2 47.1
DeiT-S + MLN (Touvron et al., 2022) Attn 58.0 M 1217.0 G 43.8 45.1
Swin-T (Liu et al., 2021) Attn 60.0 M 945.0 G 44.4 45.8
SG-Former-S (Ren et al., 2023) Attn 52.5 M 989.0 G 49.9 51.5
[…]
Mamba®-S (Wang et al., 2024j) Mamba 56.0 M - 45.3 -
PlainMamba-L2 (Yang et al., 2024a) Mamba 55.0 M 285.0 G 46.8 -
VMamba-T (Liu et al., 2024e) Mamba 62.0 M 948.0 G 48.3 48.6
FractalMamba-T (Tang et al., 2024c) Mamba 62.0 M 948.0 G 48.9 49.8
EffVMamba-S (Pei et al., 2024) Mamba 29.0 M 505.0 G 41.5 42.1
V2M-S (Wang et al., 2024m) Mamba 62 M - 48.2 49.0
GlobalMamba-T (Wang et al., 2024p) Mamba 62 M - 48.1 49.0
QuadMamba-S (Xie et al., 2024c) Mamba 62 M 961 G 47.2 48.1
MSVMamba-M (Shi et al., 2024a) Mamba 42.0 M 875.0 G 45.1 45.4
LocalVim-T (Huang et al., 2024d) Mamba 36.0 M 181.0 G 43.4 44.4
LocalVMamba-T (Huang et al., 2024d) Mamba 57.0 M 970.0 G 47.9 49.1
MambaVision-T (Hatamizadeh and Kautz, 2024) Mamba 55 M 945 G 46.6 -
GroupMamba-T (Shaker et al., 2024) Mamba 49 M 955 G 48.6 49.2
Spatial-Mamba-T (Xiao et al., 2024b) Mamba 57 M 936 G 48.6 49.4
VSSD-T (Shi et al., 2024b) Mamba 53 M 941 G 47.9 48.7
Backbone Type Params FLOPs mIoU (SS) mIoU (MS)
ConvNeXt-S (Liu et al., 2022) Conv 82.0 M 1027.0 G 48.7 49.6
MambaOut-S (Yu and Wang, 2024) Conv 76.0 M 1032.0 G 49.5 50.6
Swin-S (Liu et al., 2021) Attn 81.0 M 1039.0 G 47.6 49.5
SG-Former-M (Ren et al., 2023) Attn 68.3 M 1114.0 G 51.2 52.1
[…]
VMamba-S (Liu et al., 2024e) Mamba 82.0 M 1039.0 G 50.6 51.2
EffVMamba-B (Pei et al., 2024) Mamba 65.0 M 930.0 G 46.5 47.3
V2M-B (Wang et al., 2024m) Mamba 82 M - 50.8 51.3
GlobalMamba-S (Wang et al., 2024p) Mamba 82 M - 50.9 51.4
QuadMamba-B (Xie et al., 2024c) Mamba 82 M 1042 G 49.7 50.8
MSVMamba-T (Shi et al., 2024a) Mamba 65.0 M 942.0 G 47.6 48.5
LocalVim-S (Huang et al., 2024d) Mamba 58.0 M 297.0 G 46.4 47.5
LocalVMamba-S (Huang et al., 2024d) Mamba 81.0 M 1095.0 G 50.0 51.0
SiMBA-S (Patro and Agneeswaran, 2024b) Mamba 62.0 M 1040.0 G 49.0 49.6
MambaVision-S (Hatamizadeh and Kautz, 2024) Mamba 84 M 1135 G 48.2 -
Spatial-Mamba-S (Xiao et al., 2024b) Mamba 73 M 992 G 50.6 51.4
ConvNeXt-B (Liu et al., 2022) Conv 122.0 M 1170.0 G 49.1 49.9
MambaOut-B (Yu and Wang, 2024) Conv 112.0 M 1178.0 G 49.6 51.0
ViT-Adpt-B (Chen et al., 2023) Attn 133.9 M - 48.8 49.7
DeiT-B + MLN (Touvron et al., 2022) Attn 144.0 M 2007.0 G 45.5 47.2
Swin-B (Liu et al., 2021) Attn 121.0 M 1188.0 G 48.1 49.7
SG-Former-B (Ren et al., 2023) Attn 109.3 M 1304.0 G 52.0 52.7
[…]
Mamba®-L (Wang et al., 2024j) Mamba 377.0 M - 49.1 -
VMamba-B (Liu et al., 2024e) Mamba 122.0 M 1170.0 G 51.0 51.6
GlobalMamba-B (Wang et al., 2024p) Mamba 122 M - 51.2 51.7
MambaVision-B (Hatamizadeh and Kautz, 2024) Mamba 126 M 1342 G 49.1 -
Spatial-Mamba-B (Xiao et al., 2024b) Mamba 127 M 1176 G 51.8 52.6
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6.5em,font=,, where level=2text width=9.5em,font=,, where level=3text width=8.0em,font=,, where level=4text width=7.5em,font=,, [ Application, ver, color=carminepink!100, fill=carminepink!15, text=black [ Image (§4.1), color=lighttealblue!100, fill=lighttealblue!100, text=black, text width=5.5em [ Image Classification (§4.1.1), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. MambaTSR (Ge et al., 2024), MambaMIL (Yang et al., 2024d), Vim4Path (Nasiri-Sarvi et al., 2024), BI-Mamba (Yang et al., 2024f), MSSM (Li et al., 2024j), RSMamba (Chen et al., 2024a), color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=34em ] ] [ Image Segmentation (§4.1.2), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. HMNet (Xu et al., 2024b), ViM-UNet (Archit and Pape, 2024), Swin-UMamba (Liu et al., 2024g), VM-UNet-V2 (Zhang et al., 2024h), MM-UNet (Du et al., 2024b), Polyp-Mamba (Xu et al., 2024c),
GGVMamba (Zhou et al., 2024d), LKM-UNet (Wang et al., 2024b), SegMamba (Xing et al., 2024), EM-Net (Chang et al., 2024), CAMS-Net (Khan et al., 2025), TP-Mamba (Wang et al., 2024g), PathMamba (Fan et al., 2024),
ShapeMamba-EM (Shi et al., 2024c), Semi-Mamba-UNet (Ma and Wang, 2024), RS3Mamba (Ma et al., 2024d), ChangeMamba (Chen et al., 2024e), MambaHSI (Li et al., 2024e), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Object Detection (§4.1.3), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. FER-YOLO-Mamba (Ma et al., 2024a), VMambaCC (Ma et al., 2024e), SOAR (Verma et al., 2024), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Image Generation (§4.1.4), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. ZigMa (Hu et al., 2024), VSSM (Lambrechts et al., 2024), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Image Restoration (§4.1.5), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. FreqMamba (Zhen et al., 2024), MambaIR (Guo et al., 2024a), DVMSR (Lei et al., 2024), MambaLLIE (Weng et al., 2024), ECMamba (Dong et al., 2024c), Deform-Mamba (Ji et al., 2024),
SSUMamba (Fu et al., 2024c), FMSR (Xiao et al., 2024c), MambaFormerSR (Zhi et al., 2024), VMambaSCI (Zhang et al., 2024c), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Others (§4.1.6), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. MambaAD (He et al., 2024a), MTMamba (Lin et al., 2024), Hamba (Dong et al., 2024a), SUM (Hosseini et al., 2025), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] ] [ Video (§4.2), color=lightplum!100, fill=lightplum!100, text=black, text width=5.5em [ Video Understanding (§4.2.1), color=lightplum!100, fill=lightplum!60, text=black, text width=10em [ e.g. VideoMamba (Li et al., 2024c), VideoMamba (Park et al., 2024a), MambaTrack (Xiao et al., 2024a), DLT (Jin et al., 2024), VMRNN (Tang et al., 2024b) , color=lightplum!100, fill=lightplum!40, text=black, , text width=34em ] ] [ Video Generation (§4.2.2), color=lightplum!100, fill=lightplum!60, text=black, text width=10em [ e.g. RainMamba (Wu et al., 2024c), SSMDiff (Oshima et al., 2024), ZigMa (Hu et al., 2024), VFIMamba (Zhang et al., 2024e), MMD (Dong et al., 2024b) , color=lightplum!100, fill=lightplum!40, text=black, text width=34em ] ] ] [ Point Cloud (§4.3), color=lightblue!100, fill=lightblue!100, text=black, text width=5.5em [ e.g. PointMamba (Liang et al., 2024), Mamba3D (Han et al., 2024), 3DMambaIPF (Zhou et al., 2024c), LCM (Li et al., 2024i), VoxelMamba (Zhang et al., 2024b), MambaMOS (Zeng et al., 2024) , color=lightblue!100, fill=lightblue!40, text=black, text width=34em ] ] [ Multi-modal (§4.4), color=lightgreen!100, fill=lightgreen!100, text=black, text width=5.5em [ Homogeneous Multi-modal (§4.4.1), color=lightgreen!100, fill=lightgreen!60, text=black, text width=10em [ e.g. MGMF (Wang et al., 2024k), Sigma (Wan et al., 2025) , color=lightgreen!100, fill=lightgreen!40, text=black, text width=34em ] ] [ Heterogeneous Multi-modal (§4.4.2), color=lightgreen!100, fill=lightgreen!60, text=black, text width=10em [ e.g. InstructGIE (Meng et al., 2024), Meteor (Lee et al., 2024), RoboMamba (Liu et al., 2024d), ReMamber (Yang et al., 2024c), RSCaMa (Liu et al., 2024a), CoupledMamba (Li et al., 2024k),
MotionMamba (Zhang et al., 2024f), MambaTalk (Xu et al., 2024a), MambaGesture (Fu et al., 2024b) , color=lightgreen!100, fill=lightgreen!40, text=black, text width=34em ] ] ] ]
Figure 6. The main content flow and categorization of the application in this survey.
4. Applications
In this section, we systematically categorize and discuss the diverse vision applications of Mamba. The categorization scheme, along with an overview of the reviewed literature in this survey, is presented in Fig. 6.
4.1. Image
The processing approaches used in image applications are built upon the techniques employed in visual Mamba backbone networks, including tokenization methods, scanning strategies, and block designs, among others.
4.1.1. Classification
In addition to backbones like Vim (Zhu et al., 2024c) and VMamba (Liu et al., 2024e), which are primarily used for image classification in representation learning, the efficiency and scalability of Mamba-based architectures in handling long sequences have enabled their application across a broad range of image classification tasks (Wang et al., 2024i; Ge et al., 2024; Fang et al., 2024; Yang et al., 2024d; Nasiri-Sarvi et al., 2024; Yang et al., 2024b; Gong et al., 2024; Yue and Li, 2024; Yang et al., 2024f; Li et al., 2024j; Chen et al., 2024a). This makes them particularly effective for high-dimensional image analysis, (e.g., Whole Slide Images, 3D Medical Images, and Remote Sensing Images).
MambaTSR (Ge et al., 2024) leveraged EfficientVMamba (Pei et al., 2024) blocks to achieve efficient and accurate traffic sign recognition.
With regards to high-dimensional image analysis, MambaMIL (Yang et al., 2024d) utilized Mamba to enhance Multiple Instance Learning (MIL) for histopathology image analysis. It introduced a new approach named Sequence Reordering Mamba (SR-Mamba), which is sensitive to the order and distribution of instances and leverages the valuable information embedded within these long sequences. Differently, Vim4Path (Nasiri-Sarvi et al., 2024) utilized Vim architecture within the DINO (Caron et al., 2021) framework for representation learning. BI-Mamba (Yang et al., 2024f) employed parallel forward and backward Mamba blocks to effectively capture long-range dependencies in multi-view high-resolution chest X-rays. MSSM (Li et al., 2024j) introduced a multi-scale SSM module for 4D rs-fMRI data, specifically applied to the classification of major depressive disorder.
In the remote sensing scenario, RSMamba (Chen et al., 2024a) introduced a position-sensitive dynamic multi-path activation mechanism, encompassing forward, backward, and shuffle paths, to extract precise semantic cues for accurate scene discrimination.
Table 6. The Mamba-based methods in image modality vision tasks. The abbreviations of Data here are CXR: Chest X-Ray, CT: Computed Tomography, MRI: Magnetic Resonance Imaging, rs-fMRI: Resting-state Functional Magnetic Resonance Imaging, EM: Electron Microscopy, Echo: Echocardiograms, Micro: Microscopy Images, Skin: Skin Lesion Images, Path: Pathology. The abbreviations of Task here are CLS: Classification, DET: Detection, SEG: Segmentation, SU: Scene Understanding. Tok: Tokenization.
Methods Data Task Tok. Scan Strategy Venue Code
Direction Axis Continuity Sampling
Image Classification
MambaTSR (Ge et al., 2024) Natural Images Traffic Sign CLS 2D BD H/V Raster Global Neurocomputing24
MambaMIL (Yang et al., 2024d) Whole Slide Images Cancer Subtyping/Survival Prediction 1D SD H Raster + Reordering Global MICCAI24
Vim4Path (Nasiri-Sarvi et al., 2024) Whole Slide Images Cancer Subtyping 1D BD H Raster Global CVPR24 Workshop
BI-Mamba (Yang et al., 2024f) 2D Medical Images (X-ray) Risk Prediction 1D BD H Raster Global MICCAI24 Oral
MSSM (Li et al., 2024j) 4D Medical Images (rs-fMRI) Major Depressive Disorder CLS 1D SD H Raster Global KDD24 Workshop
RSMamba (Chen et al., 2024a) Remote Sensing Images Remote Sensing Images CLS 1D BD H/V Raster + Shuffle Global GRSL24
Image Segmentation
HMNet (Xu et al., 2024b) Natural Images Few-shot SEG 2D BD H/W Raster Global NeurIPS24
ViM-UNet (Archit and Pape, 2024) 2D Medical Images (Cell/Neurite) 2D Medical Images SEG 1D BD H Raster Global MIDL24 Short
Swin-UMamba (Liu et al., 2024g) 2D Medical Images (Endo/Micro/MRI) 2D Medical Image SEG 2D BD H/V Raster Global MICCAI24
VM-UNet-V2 (Zhang et al., 2024h) 2D Medical Images (Skin/Endo) 2D Medical Image SEG 2D BD H/V Raster Global ISBRA24
MM-UNet (Du et al., 2024b) 2D/3D Medical Images (MRI) 2D/3D Medical Image SEG 1D SD H Raster Global Scientific Reports24
Polyp-Mamba (Xu et al., 2024c) 2D Medical Images (Endo) 2D Medical Image SEG 2D BD H/V Raster Global MICCAI24
GGVMamba (Zhou et al., 2024d) 2D Medical Images (CXR) 2D Medical Image SEG 1D/2D BD H/V Raster Global MICCAI24
LKM-UNet (Wang et al., 2024b) 2D/3D Medical Images (CT/MRI) 2D/3D Medical Image SEG 1D BD H Raster Global/Local MICCAI24
SegMamba (Xing et al., 2024) 3D Medical Images (CT/MRI) 3D Medical Image SEG 1D BD + SD H + D Raster Global MICCAI24
EM-Net (Chang et al., 2024) 3D Medical Images (CT) 3D Medical Image SEG 1D SD C Raster Global MICCAI24
CAMS-Net (Khan et al., 2025) 2D Medical Images (MRI) 2D Medical Image SEG 1D BD H/C Raster Global WACV25
TP-Mamba (Wang et al., 2024g) 3D Medical Images (CT) 3D Medical Image SEG 1D SD H/V/D Raster Global MICCAI24
ShapeMamba-EM (Shi et al., 2024c) 3D Medical Images (EM) 3D Medical Image SEG 1D SD H Raster Global MICCAI24
PathMamba (Fan et al., 2024) 2D Medical Images (Path) 2D Medical Image SEG 1D BD H Raster Global MICCAI24
Semi-Mamba-UNet (Ma and Wang, 2024) 2D Medical Images (MRI) 2D Medical Image SEG 2D BD H/V Raster Global KBS24
RS3Mamba (Ma et al., 2024d) Remote Sensing Images Semantic SEG 2D BD H/V Raster Global GRSL24
ChangeMamba (Chen et al., 2024e) Remote Sensing Images Change DET/Building Damage Assessment 2D BD H/V Raster Global TGRS24
MambaHSI (Li et al., 2024e) Remote Sensing Images Remote Sensing Images CLS 1D SD H + C Raster Global TGRS24
Image Generation
ZigMa (Hu et al., 2024) Natural Images Image Generation 1D BD H/V Zigzag Global ECCV24
VSSM (Lambrechts et al., 2024) Natural Images Image Generation 1D SD H Raster Global ICML24 Workshop
Image Restoration
FreqMamba (Zhen et al., 2024) Natural Images Deraining 2D BD H/V + F Raster Global/Local MM24
MambaIR (Guo et al., 2024a) Natural Images Super-resolution/Denoising 2D BD H/V Raster Global ECCV24
DVMSR (Lei et al., 2024) Natural Images Super-resolution 1D SD H Raster Global CVPR24 Workshop
MambaLLIE (Weng et al., 2024) Natural Images Enhancement 2D BD H/V Raster Global NeurIPS24
ECMamba (Dong et al., 2024c) Natural Images Enhancement 2D BD H/V Reordering Global NeurIPS24
Deform-Mamba (Ji et al., 2024) 2D Medical Images (MRI) Super-resolution 2D BD H/V Raster Global MICCAI24
SSUMamba (Fu et al., 2024c) Remote Sensing Images Hyperspectral Denoising 1D BD H/V/D Raster Global TGRS24
FMSR (Xiao et al., 2024c) Remote Sensing Images Super-resolution 2D BD H/V Raster Global TMM24
MambaFormerSR (Zhi et al., 2024) Remote Sensing Images Super-resolution 2D BD H/V Raster Global GRSL24
VMambaSCI (Zhang et al., 2024c) Remote Sensing Images Spectral Reconstruction 2D BD H/V + C Raster Global MM24
Others
MambaAD (He et al., 2024a) Natural Images Multi-class Anomaly DET 2D BD H/V/H*/V* Hilbert Global NeurIPS24
MTMamba (Lin et al., 2024) Natural Images Multi-task Dense SU 2D BD H/V Raster Global ECCV24
Hamba (Dong et al., 2024a) Natural Images 3D Hand Reconstruction 2D BD Graph-guided Raster Global NeurIPS24
SUM (Hosseini et al., 2025) Natural Images Visual Attention Modeling 2D BD H/V Raster Global WACV25
4.1.2. Segmentation
Segmentation remains a vital and prominent area in computer vision, holding immense value for diverse real-world applications (Chen et al., 2024f; Du et al., 2024b; Ma et al., 2024b; Xing et al., 2024; Ruan and Xiang, 2024; Liu et al., 2024g; Gong et al., 2024; Wang et al., 2024o; Ma and Wang, 2024; Ye and Chen, 2024; Wang and Ma, 2024; Liao et al., 2024; Wang et al., 2024a; Zhang et al., 2024h; Wu et al., 2024a; Wang et al., 2024b; Xie et al., 2024b; Tang et al., 2024a; Sanjid et al., 2024a; Wu et al., 2024b; Hao et al., 2024; Archit and Pape, 2024; Isensee et al., 2024; Sanjid et al., 2024b; Nguyen et al., 2024; Xu, 2024; Yuan et al., 2024; Tsai et al., 2024; Chen et al., 2024d; Shi et al., 2024c; Wang et al., 2024g; Chang et al., 2024; Fan et al., 2024; Zhou et al., 2024d; Zhu et al., 2024a; Ma et al., 2024d; Yao et al., 2024; Wang et al., 2024l; Huang et al., 2024a; Zhu et al., 2024b; Liu et al., 2024b; Zhao et al., 2024a; Chen et al., 2024e; Zhou et al., 2024a; He et al., 2024c; Li et al., 2024e; Malik et al., 2024; Wei et al., 2024; Khan et al., 2025). Mamba is well-positioned to enhance segmentation tasks for its strong long-range modeling capability.
In the context of few-shot segmentation, HMNet (Xu et al., 2024b) identified two issues when concatenating support and query features and applying the original Mamba: support forgetting and the intra-class gap between support and query. To mitigate these issues, HMNet introduced two mechanisms: periodically re-scanning the support features during the query feature scanning process, and scanning the query features in parallel to avoid undesired pixel-level interactions within the query features.
Recent advances based on Mamba have appeared in the medical domain, encompassing diverse 2D and 3D medical image datasets (Tsai et al., 2024; Chen et al., 2024d). Several methods (Archit and Pape, 2024; Liu et al., 2024g; Zhang et al., 2024h) directly replace CNN-blocks in the U-Net architecture with Mamba-based blocks (Gu and Dao, 2024; Liu et al., 2024e; Zhu et al., 2024c). U-Mamba (Ma et al., 2024b) is the first effective attempt to apply Mamba in medical image segmentation. It introduced a hybrid CNN-SSM block. ViM-UNet (Archit and Pape, 2024) utilized Vim (Zhu et al., 2024c) block as the basic blocks. Swin-UMamba (Liu et al., 2024g) and VM-UNet-V2 (Zhang et al., 2024h) employed the VMamba (Liu et al., 2024e) block as a basic block to capture extensive contextual information. MM-UNet (Du et al., 2024b), Polyp-Mamba (Xu et al., 2024c), GGVMamba (Zhou et al., 2024d), and LKM-UNet (Wang et al., 2024b) employed Mamba-based architecture as a basic block with additional designs for enhancement. SegMamba (Xing et al., 2024) introduced gated spatial convolution and tri-orientated Spatial Mamba, which utilized an inter-slice path for 3D data. EM-Net (Chang et al., 2024) employed a channel scanning strategy and integrated frequency analysis. CAMS-Net (Khan et al., 2025) introduced Mamba-based channel and spatial aggregators. TP-Mamba (Wang et al., 2024g) and ShapeMamba-EM (Shi et al., 2024c) each introduced a 3D adapter based on Mamba to fine-tune the foundation models SAM (Kirillov et al., 2023) and SAM-Med3D (Wang et al., 2024d), respectively, for medical image data. PathMamba (Fan et al., 2024) incorporated Mamba into weakly supervised multi-label segmentation of pathology images and introduced a dual-granularity comparative Mamba block, designed to facilitate pixel- and patch-level contrastive learning. In (Ma and Wang, 2024), a purely Mamba-based UNet and a CNN-based UNet were integrated into a semi-supervised learning framework, where they perform cross-supervision and contrastive learning.
In the field of remote sensing images, Mamba-based approaches typically leverage its efficiency for processing high-resolution images (Ma et al., 2024d), and exploit the spatial-temporal and spatial-spectral characteristics of images to perform scanning (Chen et al., 2024e; Li et al., 2024e).
4.1.3. Detection
The advantages of Mamba in extracting global information and performing efficient computations make it particularly well-suited for object detection tasks.
FER-YOLO-Mamba (Ma et al., 2024a) is the first Mamba-based model for facial expression detection. It introduced a dual-branch structure that incorporated omnidirectional Mamba and attention mechanisms for the feature pyramid network. VMambaCC (Ma et al., 2024e) is a pioneering work that applied the VMamba (Liu et al., 2024e) architecture to crowd counting tasks. SOAR (Verma et al., 2024) integrated vim (Zhu et al., 2024c) block into the YOLO v9 architecture.
4.1.4. Generation
Intuitively, applying the Mamba architecture to a series of generation tasks to achieve sufficient long sequence interactions has the potential to achieve impressive performance (Fei et al., 2024; Hu et al., 2024; Teng et al., 2024; Mo and Tian, 2024; Lambrechts et al., 2024; Shen et al., 2024; Fu et al., 2024a; Ju and Zhou, 2024; Qian et al., 2024).
ZigMa (Hu et al., 2024) is a DiT-style (Peebles and Xie, 2023) Mamba diffusion model. It leveraged spatial continuity to maximally incorporate the inductive bias from visual data. VSSM (Lambrechts et al., 2024) is a variational autoencoder in which both the encoder and decoder are SSMs, enabling efficient parallel computation.
4.1.5. Restoration
Recently, the Mamba architecture has also been extensively applied to various image restoration tasks. These low-level tasks include reconstructing degraded images to their original state, such as image dehazing (Zheng and Wu, 2024; Zhou et al., 2024b), deraining (Zheng and Wu, 2024; Zhen et al., 2024; Yamashita and Ikehara, 2024; Li et al., 2024d; Shi et al., 2024d), denoising (Deng and Gu, 2024; Guo et al., 2024a; Liu et al., 2024f; Fu et al., 2024c), deblurring (Deng and Gu, 2024; Gao and Dang, 2024; Kong et al., 2024), super-resolution (Guo et al., 2024a; Shi et al., 2024d; Cheng et al., 2024; Lei et al., 2024; Huang et al., 2024c; Ji et al., 2024; Xiao et al., 2024c; Zhi et al., 2024), and other reconstruction (Zheng and Zhang, 2024; Huang et al., 2024g, h; Li et al., 2024g; Wang et al., 2024e; Zhang et al., 2024c), and enhancing the image quality by adjusting luminance, contrast, and other visual attributes, such as multiple exposure correction (Tan et al., 2024), low light enhancement (Zheng and Wu, 2024; Weng et al., 2024; Bai et al., 2024), and underwater image enhancement (Chen and Ge, 2024; Guan et al., 2024).
The extraction of local fine-grained features plays a critical role in tackling low-level image tasks, which the vanilla Mamba architecture lacks. Besides, image restoration approaches often utilize techniques from the frequency domain, leveraging the ability to isolate and manipulate specific image frequency components for improving detail recovery. For instance, FreqMamba (Zhen et al., 2024) utilized both Global and Local scanning, effectively removing high-frequency rain streaks in deraining tasks by leveraging frequency-based techniques. Similarly, in image super-resolution tasks, MambaIR (Guo et al., 2024a) incorporated convolutions and channel attentions to enhance the Mamba architecture’s capacity for preserving local information while effectively reducing channel redundancy. Differently, DVMSR (Lei et al., 2024) applied a distillation strategy to the Mamba-based network to enhance its efficiency. In image enhancement, MambaLLIE (Weng et al., 2024) enhanced local dependencies by incorporating an augmented local bias within the 2D selective scan mechanism and introducing an implicit Retinex-aware selective kernel module. ECMamba (Dong et al., 2024c) integrated Retinex theory for exposure correction and presented a feature-aware scanning strategy that leverages an activation response map from deformable convolution to generate sequences ordered by feature importance.
In the medical domain, Deform-Mamba (Ji et al., 2024) combined the modulated deformable block with VMamba (Liu et al., 2024e) block to effectively capture both local and global information. Additionally, it introduced a multi-view context module and a contrastive edge loss.
In the remote sensing domain, to improve computational efficiency in hyperspectral denoising, SSUMamba (Fu et al., 2024c) introduced a spatial-spectral continuous scan Mamba, alternating rows, columns, and bands in six different orders to capture long-range spatial-spectral dependencies. Furthermore, 3D convolutions are incorporated to enhance local spatial-spectral modeling. In remote sensing image super-resolution tasks, FMSR (Xiao et al., 2024c) enhanced spatial-frequency fusion by integrating frequency selection modules with the global VMamba (Liu et al., 2024e) block and a local hybrid gating module. MambaFormerSR (Zhi et al., 2024) combined Mamba with Transformer architectures, while also introducing a convolutional Fourier transform feed-forward network. VMambaSCI (Zhang et al., 2024c) introduced a novel spatial-channel scanning technique for spectral reconstruction.
4.1.6. Others
Mamba’s ability to model long-range dependencies extends its advantages to various other image modality tasks (He et al., 2024a; Qin et al., 2024; Lin et al., 2024; Dong et al., 2024a; Hosseini et al., 2025), improving feature representation and contextual understanding across diverse applications. MambaAD (He et al., 2024a) applied Mamba to multi-class unsupervised anomaly detection by combining a pre-trained encoder with a Mamba-based decoder. The decoder integrated parallel cascaded hybrid state space blocks, utilizing Hilbert scanning in eight directions, and multi-kernel convolution operations at multiple scales to effectively capture both global and local information. MTMamba (Lin et al., 2024) introduced self-task and cross-task Mamba blocks for multi-task dense prediction. In particular, the cross-task Mamba block was designed to enhance interactions between tasks. Hamba (Dong et al., 2024a) introduced a graph-guided bidirectional scanning technique for 3D hand reconstruction from a single RGB image, efficiently utilizing a few effective tokens to learn the joint relations and spatial sequences. SUM (Hosseini et al., 2025) leveraged Mamba for efficient long-range dependency modeling and introduced a novel conditional visual state space block, enabling dynamic adaptation to different image types, in visual attention modeling tasks.
Table 7. The Mamba-based methods in video modality vision tasks. The abbreviations of Task here are CLS: Classification, VU: Video Understanding, T2VR: Text-to-Video Retrieval, MOT: Multi-object Tracking, STF: Spatiotemporal Forecasting, VFI: Video Frame Interpolation.
Methods Data Task Scan Strategy Venue Code
Video Understanding
VideoMamba (Li et al., 2024c) Natural Videos Action CLS, VU, T2VR BD Spatial-first ECCV24
VideoMamba (Park et al., 2024a) Natural Videos VU BD Spatial-first ECCV24
MambaTrack (Xiao et al., 2024a) Natural Videos MOT BD Spatial-first MM24 Oral
DLT (Jin et al., 2024) Natural Videos MOT BD Spatial-first MM24
VMRNN (Tang et al., 2024b) Natural Videos STF Spatial CVPR24 Workshop
Video Generation
RainMamba (Wu et al., 2024c) Natural Videos Video Deraining BD Spatial-first (Zigzag) + 3D Hilbert MM24 Oral
SSMDiff (Oshima et al., 2024) Natural Videos Video Generation Spatial + Temporal ICLR24 Workshop
ZigMa (Hu et al., 2024) Natural Videos Video Generation Spatial + Temporal ECCV24
VFIMamba (Zhang et al., 2024e) Natural Videos VFI Temporal-first NeurIPS24
MMD (Dong et al., 2024b) Motion Sequence Motion Generation Spatial + Temporal MM24
Figure 7. Video scanning techniques.
[…]
4.2. Video
Video processing is a key area in computer vision, dealing with 3D video data that consists of sequences of 2D frames. Video data includes a temporal dimension in addition to the spatial dimensions, enabling the capture of temporal dynamics. The primary goal of video processing is to effectively master spatio-temporal representations across long contexts. Mamba excels in this domain with its selective state space model, which achieves a balance between maintaining linear complexity and enabling effective long-term dynamic modeling. To adapt Mamba to 3D video data, current approaches organize the 3D video tokens in the following ways.
Spatial-first scans spatial tokens by location and then stacks them frame by frame. The scanning techniques for spatial tokens within each frame are similar to those used for 2D images. Temporal-first arranges temporal tokens by frame and then stacks them along the spatial dimension. To improve the receptive field and mitigate Mamba’s position sensitivity, these spatiotemporal scanning paths can be directly reversed. Considering the sequential attribute of the temporal dimension, Spatial Reversal reverses the spatial tokens in each frame and maintains the order along the temporal axis. Temporal Reversal maintains the order of spatial tokens and reverses the temporal sequence.
Additionally, some approaches apply Mamba separately to spatial and temporal information, utilizing 2D Spatial scanning and 1D Temporal scanning. Other approaches focus solely on Spatial scanning, leveraging alternative techniques to handle temporal dynamics. The video data can also be regarded as 3D data. In this situation, scanning techniques like 3D zigzag scanning and 3D Hilbert scanning can be utilized.
These innovative approaches, illustrated in Fig. 7, have been widely adopted in diverse video analysis tasks such as video understanding (Li et al., 2024c) and video generation (Hu et al., 2024).
4.2.1. Video Understanding
Recent approaches in video understanding tasks have adopted Mamba as the core architecture due to its superior long-range modeling capability and computational efficiency for analyzing high-resolution long video data, with a primary focus on optimizing scanning techniques (Yang et al., 2024e; Li et al., 2024c; Park et al., 2024a; Chen et al., 2024c; Zou et al., 2024b; Xiao et al., 2024a; Jin et al., 2024; Tang et al., 2024b; Chen et al., 2024b; Liu et al., 2024c). Both VideoMamba (Li et al., 2024c) and VideoMamba (Park et al., 2024a) built a purely Mamba-based model by stacking identical bi-directional Mamba blocks and utilizing the BD Spatial-first scanning strategy. VideoMamba (Li et al., 2024c) conducted extensive experiments comparing Spatial-first scanning, Temporal-first scanning, and their combination. The results indicate that BD Spatial-first scanning achieves the best performance, highlighting VideoMamba’s strong scalability, sensitivity to short-term action recognition, superior long-term video understanding, and compatibility with other modalities. VideoMamba (Park et al., 2024a) explored the reverse of Spatial-first scanning, conducting experiments on direct backward scanning, Spatial Reversal, and Temporal Reversal. The results reaffirm that BD Spatial-first scanning yields the best performance. Similarly, the Mamba architecture is employed for 3D long-term modeling in the multi-object tracking paradigm. MambaTrack (Xiao et al., 2024a) leveraged Mamba to model the spatio-temporal dynamics of objects and predict their future motion. DLT (Jin et al., 2024) employed Mamba to process multiple frames and estimate the distances of pre-detected 2D objects in these frames from a monocular camera. VMRNN (Tang et al., 2024b) integrated the 2D visual Mamba block with LSTM for spatiotemporal forecasting.
4.2.2. Video Generation
Video generation approaches focus on capturing correlations across the spatiotemporal dimension (Wu et al., 2024c; Oshima et al., 2024; Hu et al., 2024; Gao et al., 2024; Mo and Tian, 2024; Zhang et al., 2024e; Dong et al., 2024b). RainMamba (Wu et al., 2024c) introduced a 3D Hilbert scanning mechanism, while SSMDiff (Oshima et al., 2024), Zigma (Hu et al., 2024), and MMD (Dong et al., 2024b) decoupled spatial and temporal modeling. VFIMamba (Zhang et al., 2024e) utilized Temporal-first scanning to enhance spatiotemporal adjacency for the frame interpolation task.
Table 8. The Mamba-based methods in point cloud vision tasks. The abbreviations of Task here are CLS: Classification, DET: Detection, SEG: Segmentation. Tok: Tokenization.
Methods Data Task Tok. Scan Strategy Venue Code
PointMamba (Liang et al., 2024) 3D Point Cloud Data Object CLS, Part SEG FPS & KNN & Hilbert/Trans-Hilbert SD Raster NeurIPS24
Mamba3D (Han et al., 2024) 3D Point Cloud Data Object CLS, Part SEG FPS & KNN Raster + Channel Flipping MM24
LCM (Zha et al., 2024) 3D Point Cloud Data Object CLS, Object DET, Part SEG FPS & KNN SD Raster NeurIPS24
VoxelMamba (Zhang et al., 2024b) 3D Point Cloud Data Object DET Voxelization Multi-scale BD Hilbert NeurIPS24
MambaMOS (Zeng et al., 2024) 4D Point Cloud Data Object SEG Z/Hilbert SD Raster MM24
4.3. Point Cloud
Point clouds represent a fundamental 3D data structure, consisting of discrete points in space, each defined by a set of 3D coordinates. These points collectively describe the geometry of objects or scenes. However, the inherent sparse and irregular distribution of point clouds poses significant challenges for 3D vision tasks. Point clouds can be used directly as raw points or transformed into regular grids through voxelization for further processing. When using raw points, they must first undergo a serialization process during tokenization. Notably, the computational complexity of 3D point cloud analysis can be reduced by grouping points or voxels. A common approach for grouping raw points during tokenization involves sampling key points using Farthest Point Sampling (FPS), followed by selecting k-nearest neighbors (KNN) for each key point.
Inspired by the linear complexity and global modeling capabilities of Mamba, several general SSM-based backbones (Zubic et al., 2024; Liang et al., 2024; Zhang et al., 2024d; Liu et al., 2024h; Han et al., 2024; Wang et al., 2024c; Zha et al., 2024; Zhang et al., 2024b; Zhou et al., 2024c; Li et al., 2024i; Ren et al., 2024b; Xiang et al., 2024; Zeng et al., 2024; Liu et al., 2024c; Mo, 2024) have been explored in point cloud processing. Adaptations of 1D causal Mamba operations for 3D point clouds typically focus on scanning strategies and tokenization of raw points. For example, Z-order and Hilbert curves can be employed during scanning or tokenization to preserve spatial locality.
PointMamba (Liang et al., 2024) applied Hilbert and Trans-Hilbert curves to the key points obtained through FPS before performing KNN during tokenization. This study demonstrated that PointMamba, leveraging a simple Mamba block, outperforms state-of-the-art Transformer-based networks. Mamba3D (Han et al., 2024) identified that learning sequence order in point clouds can lead to unreliable and unstable pseudo-order dependencies. To mitigate this, channel flipping was introduced for the backward SSM. LCM (Zha et al., 2024) applied the Mamba architecture to masked point modeling for pre-training, and proposed a locally constrained compact point cloud model. VoxelMamba (Zhang et al., 2024b) introduced a Multi-scale BD Hilbert scanning strategy. MambaMOS (Zeng et al., 2024) employed Z-order and Hilbert curves to serialize 4D points and designed a new embedding layer to better capture spatial and temporal information.
Table 9. The Mamba-based methods in Multi-modal Vision Tasks. The abbreviations of Task here are DET: Detection, SEG: Segmentation, SU: Scene Understanding, VQA: Visual Question Answering. The abbreviations of Block here are VSS: Vision State Space Module, *: modification.
Methods Data Task Fusion Strategy Venue Code
Paradigm Mamba Location Mamba Design
Homogeneous Multi-modal
MGMF (Wang et al., 2024k) RGB Images & Infrared Images Object DET Multi-branch Fusion Module M TGRS24
Sigma (Wan et al., 2025) RGB Images & Thermal/Depth Images Semantic SEG Multi-branch Backbone/Fusion Module/Decoder VSS/M/U/VSS* WACV25
Heterogeneous Multi-modal
InstructGIE (Meng et al., 2024) Natural Images & Text Image Editing Diffusion Vision Encoder VSS ECCV24
Meteor (Lee et al., 2024) Natural Images & Text VQA, Reasoning, SU MLLM LLM Mamba NeurIPS24
RoboMamba (Liu et al., 2024d) Natural Images & Text Robot Reasoning and Manipulation MLLM LLM Mamba NeurIPS24
ReMamber (Yang et al., 2024c) Natural Images & Text Referring Image SEG Uni-branch Backbone U ECCV24
RSCaMa (Liu et al., 2024a) Remote Sensing Images & Text Change Captioning Multi-branch Fusion Module M/U GRSL24
CoupledMamba (Li et al., 2024k) Audio & Video & Text Sentiment Analysis Uni-branch Backbone M NeurIPS
MotionMamba (Zhang et al., 2024f) Motion Sequence & Text Motion Generation Diffusion Backbone CA ECCV24
MambaTalk (Xu et al., 2024a) Audio & Text & Motion Motion Generation VQ-VAE Fusion Module CA/U NeurIPS24
MambaGesture (Fu et al., 2024b) Audio & Text & Style & Emotion Motion Generation Diffusion Backbone CA/A MM24
4.4. Multi-modal
Multi-modal tasks have become increasingly pivotal in the domain of computer vision, enabling the enrichment of visual understanding through the integration of diverse information sources. The fundamental objective of multi-modal tasks is to learn valuable potential feature representations from a multitude of modalities, such as textual captions and visual images, RGB images with supplementary components like depth or thermal images, and various forms of medical imaging data. Nevertheless, a significant challenge in attaining multi-modal objectives lies in the effective capture and amalgamation of the information across diverse modalities.
Early multi-modal approaches employ the Multi-branch paradigm, where each branch processes distinct modalities in parallel, followed by the fusion of features to create a unified representation. With the advent of Transformer architectures, a new Uni-branch paradigm has emerged, which concatenates various modalities and processes them within a single branch, simplifying the integration process and achieving greater efficiency. To harness the potential of large language models (LLMs), the Multi-modal Large Language Model (MLLM) paradigm utilizes LLMs as task decoders, facilitating enhanced multi-modal understanding. Additionally, established paradigms such as Diffusion (Ho et al., 2020) and VQ-VAE (van den Oord et al., 2017) are widely adopted for multi-modal generation tasks, leveraging their robust generative capabilities.
Recently, Mamba has been adapted for a variety of multi-modal tasks due to its computational efficiency and flexibility in feature fusion, demonstrating impressive performance across different applications. Firstly, various modules within existing paradigms can be replaced with Mamba-based architectures to leverage its global receptive capabilities and linear computational complexity, such as backbones and vision encoders. Notably, the LLM component of MLLM paradigms can be substituted with Mamba to capitalize on its advantages. Secondly, leveraging Mamba’s selective SSM mechanism, novel Mamba designs can enhance feature fusion, enabling flexible interactions between features from different modalities. Some fusion approaches employ distinct SSM mechanisms for each modality, integrating information from other modalities (individual, combined, or differences between modalities) into the SSM framework (M). Some fusion approaches concatenate information from different modalities, leveraging the SSM mechanism to fuse them through advanced scanning techniques (U). Additionally, mechanisms such as Cross Attention (CA) and Attention (A) can be incorporated into the Mamba-based architecture to facilitate effective feature fusion.
[…]
4.4.1. Homogeneous Multi-modal
Homogeneous multi-modal applications refer to the tasks where the input data consists of multiple modalities sharing similar data type, including MRI-CT registration (Guo et al., 2024c; Wang et al., 2024n), medical image fusion (Li et al., 2024f; Xie et al., 2024a), medical image generation (Atli et al., 2024), remote sensing images pansharpening (He et al., 2024b; Peng et al., 2024; Cao et al., 2024), infrared-visible image fusion (Li et al., 2024f; Xie et al., 2024a; Ma et al., 2024c), RGB-infrared image object detection (Dong et al., 2024d; Li et al., 2024b, f; Wang et al., 2024k) and RGB-thermal/depth image semantic segmentation (Wan et al., 2025).
For RGB-infrared object detection, the multi-branch paradigm MGMF (Wang et al., 2024k) introduced the Mamba-based fusion module, while utilizing the common feature extraction backbone and detection head. In particular, this fusion module maps the summed features from the two modalities into a hidden state space and employs gating mechanisms to facilitate state transitions, enabling deep cross-modal fusion of the representations. Sigma (Wan et al., 2025) is another multi-branch paradigm designed for multi-modal semantic segmentation, employing a fully Mamba-based architecture across all components. Specifically, VMamba (Liu et al., 2024e) blocks are used in the feature extraction backbone. For the fusion module, cross-modal multiplication and feature concatenation mechanisms are integrated within the SSM. Furthermore, channel attention is incorporated into the SSM to enhance the decoder’s capability in modeling channel-wise interactions.
4.4.2. Heterogeneous Multi-modal
Heterogeneous multi-modal applications refer to tasks where the input data consists of multiple modalities with different data types (Meng et al., 2024; Qiao et al., 2024; Lee et al., 2024; Zhao et al., 2024b; Liu et al., 2024d; Yang et al., 2024c; Huang et al., 2024e; Chen et al., 2024g; Liu et al., 2024a; Li et al., 2024a; Shou et al., 2024; Li et al., 2024k; Huang et al., 2024f; Zhang et al., 2024f; Fu et al., 2024b; Wang et al., 2024f; Sun et al., 2024).
In the field of text-driven generation, Mamba architectures are utilized for visual processing. To enhance the generalization capabilities of image editing approaches, InstructGIE (Meng et al., 2024) reformed the conditional latent diffusion model by integrating a vision encoder based on VMamba (Liu et al., 2024e), and used the VMamba layer to inject the processed visual condition information into the frozen Stable Diffusion (Rombach et al., 2022) model. In the realm of MLLMs, Meteor (Lee et al., 2024) introduced multifaceted rationales and applied the Mamba architecture to embed these lengthy rationales to enhance the understanding and answering capabilities of MLLMs. RoboMamba (Liu et al., 2024d) aligned the vision encoder with an efficient Mamba language model to enable cost-effective robot manipulation capabilities. ReMamber (Yang et al., 2024c) introduced a novel referring image segmentation framework containing multiple Mamba twister blocks. Each block comprises VMamba (Liu et al., 2024e) layers to extract the vision feature and a twister layer to integrate the textual information into the visual modality. The Twister layer first constructs a hybrid feature cube and subsequently performs channel-wise and spatial scanning for feature fusion. RSCaMa (Liu et al., 2024a) employed multiple CaMa layers between the backbone and the language decoder to facilitate efficient joint spatial-temporal modeling for remote sensing image change captioning. Specifically, the CaMa layer enhances change awareness by multiplying the differencing features with the output of SSMs. It further employs a Temporal-first scanning strategy to facilitate effective temporal modeling.
CoupledMamba (Li et al., 2024k) proposed coupling state chains of inter-modalities while maintaining the independence of intra-modality state processes. It also derived the global convolution kernel to enable parallelism for the coupled state transition scheme. MotionMamba (Zhang et al., 2024f) incorporated two modules based on the vanilla Mamba into a diffusion-based generative system for text-to-motion synthesis, namely hierarchical temporal Mamba and bidirectional spatial Mamba. MambaGesture (Fu et al., 2024b) integrated Mamba architecture with the attention mechanism to form the basic block for the diffusion model in the co-speech gesture generation task.
5. Challenges and Future Directions
In this section, we identify current challenges and outline promising directions, to enhance Mamba’s capabilities and broaden its applications in computer vision.
5.1. Scalability
5.1.1. Model Size
In the current era of large foundation models, the scalability of deep neural networks is pivotal for advancing general artificial intelligence. Mamba is emerging as a promising foundational architecture, offering an alternative to prevalent Transformers with its ability to scale linearly with sequence length. However, while existing techniques aimed at reducing the computational complexity of Transformers often compromise on model expressivity or adaptability, Mamba also encounters challenges when scaled to large-sized networks.
Firstly, the Mamba architecture faces stability issues when scaled to large network configurations (Patro and Agneeswaran, 2024a). The underlying causes of Mamba’s instability remain inadequately understood. This instability often manifests as vanishing or exploding gradients (Patro and Agneeswaran, 2024b), leading to performance degradation or even training collapse. To address the stability issues, some studies utilize stable reparameterization to balance gradient scales in state space models (Wang and Li, 2024; Ran-Milo et al., 2024; Orvieto et al., 2024). Some works utilize learnable Fourier Transforms to adjust the eigenvalues of the evolution matrix, ensuring they are negative real numbers, which helps maintain stability (Patro and Agneeswaran, 2024b). Mamba-2 incorporates an additional normalization layer in the block before the final output projection, mitigating instability issues encountered in large models (Dao and Gu, 2024a). Despite these efforts, most visual Mamba models are limited to base or smaller scales, which may restrict their performance.
Secondly, visual Mamba models, when controlled for similar parameters or FLOPs, are generally outperformed by cutting-edge CNNs, advanced visual Transformers, and hybrid architectures that integrate multiple model types. This performance gap also hinders the scaling of visual Mamba models and underscores the necessity to enhance their performance. Integrating architectures or techniques proven effective in CNNs and visual Transformers could help bridge the performance gap. Besides, while there may be some compromise in performance, the non-hierarchical architectures comprising identical blocks facilitate optimization for hardware acceleration. This aspect is crucial for scaling to large models and enhancing computational efficiency. In addition, self-supervised pre-training plays a critical role in enhancing the model performance. For instance, ARM (Ren et al., 2024a) demonstrated that the autoregressive pre-training can significantly improve the performance of the Mamba architecture and address the instability issues, efficiently unlocking its potential for scaling to very large model sizes. However, the self-supervised pre-training strategies for Mamba models are still nascent.
More recently, Mamba-2 (Dao and Gu, 2024a) outlined several potential directions for further enhancing its performance. Given that SSD enhances hardware efficiency by limiting expressivity through the use of restricted structured matrices, future work could explore refining structured matrix algorithms toward more generalized diagonal SSMs to balance efficiency and expressivity (Merrill et al., 2024). Mamba-2 further suggests that advancements from a matrix mixer perspective could broaden the applications of sequence models, including the development of principled non-causal Mamba variants and methods to bridge the gap between softmax attention and sub-quadratic models by analyzing their matrix transformation structures.
Future research should explore optimizing these aspects to harness the full capabilities of large network configurations.
5.1.2. Data
The Mamba architecture theoretically reduces computational complexity, enhancing its efficiency in long sequence processing. Moreover, unlike many sequence models that do not benefit from extended contexts (Shi et al., 2023), Mamba can selectively disregard irrelevant information, yielding consistent performance improvements as context length increases. The efficiency and efficacy of Mamba architecture bring significant opportunities for data scalability.
Firstly, the Mamba architecture offers significant potential for effectively handling high-dimensional data by treating each sample as a long sequence. This capability is particularly beneficial for processing high-resolution and multi-dimensional data including high-resolution remote sensing images, whole slice pathology images, highspectral and multispectral images, 3D medical images, and point clouds. Secondly, its efficiency and prowess in long sequence modeling not only facilitate scaling to large-scale datasets, but also enable effective analysis of the temporal or spatial relationships among samples. The typical datasets include multi-temporal remote sensing images, where changes over time provide essential information, and long-term video frames, where continuous activities must be analyzed over extended periods.
Furthermore, given the parameter efficiency of the Mamba models and their inherent local inductive bias, which typically mitigate overfitting, Mamba models have considerable potential for achieving optimal performance without relying on large-scale datasets. This opens up opportunities for developing small yet effective Mamba-based models advantageous in resource-constrained environments, a pathway that is underexplored in current research.
Future research should concentrate on scaling Mamba models for processing visual data with higher dimensions and larger scales. Fully exploiting the visual data by designing small models that maintain high performance is also a promising future direction.
5.1.3. Hardware-aware
Previous SSMs such as S4 are time- and input-invariant, making it feasible to unroll the recurrent representation into the convolutional representation for efficient parallel computation during training. However, the selection mechanism in Mamba makes the model become time- and input-dependent, making it infeasible to train the model under the convolutional representation. To overcome this limitation of the selective SSMs (Mamba) and make them efficient on modern hardware, hardware-aware algorithm has been designed to accelerate the training under the recurrent mode.
Mamba-1 uses kernel fusion and recomputation to make the computation of the parallel SSM scan fast and memory-efficient. Specifically, the discretization step, the scan, and the multiplication with the projection matrix are fused into one kernel. Then the SSM parameters are loaded from slow HBM to fast SRAM for calculating the fused kernel in SRAM. Subsequently, the outputs are written back to HBM. In addition, Mamba-1 chooses to recompute the intermediate states for the backpropagation since the recomputation is faster than storing them and reading them from HBM. These two strategies accelerate the scan operation and reduce the GPU requirement. However, according to research findings (Wang et al., 2024h; Dao and Gu, 2024a), GPU consumption for the Mamba model does not consistently demonstrate a reduction compared to the Transformer model, highlighting the need for optimized, hardware-aware Mamba algorithms tailored for vision tasks.
Mamba-2 (Dao and Gu, 2024a) establishes theoretical connections between structured SSMs and variants of attention by reformulating different approaches to computing SSMs as various matrix multiplication algorithms on structured matrices and generalizing the linear attention (Katharopoulos et al., 2020) to structured masked attention. Based on this theoretical analysis, Mamba-2 introduces an efficient SSD algorithm for SSM computation by employing block decompositions of particular semiseparable matrices and modifies the Mamba block to be tensor parallelism friendly. This framework allows Mamba-2 to leverage system optimizations developed for Transformers, including tensor parallelism and sequence parallelism, facilitating the parallel training of large models using multiple GPUs. Moreover, unlike Transformers, Mamba-2 can be efficiently trained with variable sequence lengths. Given that semiseparable matrices are extensively studied in scientific computing, incorporating these techniques presents promising opportunities for further advancing SSMs (Dao and Gu, 2024a). Besides, the adaptation of Mamba-2 to vision tasks requires further investigation.
Future research should focus on developing hardware-aware approaches to enhance the computational efficiency of visual Mamba models. On the one hand, hardware-aware approaches that reduce computational overhead are essential for scaling to larger hardware configurations. On the other hand, from a broader perspective, they are crucial for the development of green artificial intelligence technologies, which attract more and more attention from the community due to the need for energy consumption reduction to protect the environment.
5.1.4. Task
The Mamba architecture, noted for its efficiency and proficiency in long sequence modeling, enables applications across a broader spectrum of tasks compared to CNNs and Transformers.
Firstly, as previously mentioned, the Mamba architecture offers broader possibilities for handling high-dimensional data and large-scale datasets.
Secondly, the Mamba model’s proficiency in handling extended sequences significantly enhances its applicability in multi-modal learning. Similar to the Transformer, which effectively models both natural languages and images within a unified framework, the long sequence modeling capability of the Mamba architecture allows it to involve more modalities such as time series and extended textual content. Besides, like the Transformer, Mamba can adeptly handle concatenated independent sequences, processing them either as a whole or as distinct entities by resetting their states at boundaries. Furthermore, Mamba offers computational efficiency and time-scale robustness, the latter of which is particularly advantageous for addressing issues with missing modalities in data streams.
Thirdly, the high efficiency of the Mamba architecture enables the development of high-speed and highly efficient applications. This includes interactive systems, autonomous driving, and surveillance, where real-time processing and reduced computational overhead are critical.
Future research should focus on effectively leveraging the Mamba architecture to enhance performance in existing tasks and to explore new applications.
5.2. Causality
In visual Transformer-based architectures, the attention matrix explicitly captures dependencies between each pair of visual tokens, enabling predictions at any location to have a global receptive field. However, in Mamba-based architectures, the S6 blocks make predictions for each token based solely on the hidden state $h_{t-1}$ and the current input $x_{t}$. The $h_{t-1}$ only incorporates information from previously scanned tokens and disregards information from other tokens. Consequently, the vanilla Mamba is a causal system where predictions rely solely on the current and previous inputs. Although the causality is naturally suitable for autoregressive models in sequential vision tasks (Ren et al., 2024a), it contradicts the nature of spatial images, resulting in a fundamental mismatch between Mamba-based architectures and common visual tasks. To address this issue, existing works have proposed to integrate multiple scanning techniques to ensure a global receptive field. While these strategies partially mitigate the undesirable causality of Mamba in vision tasks, they still encounter challenges or introduce new drawbacks that require resolution.
Firstly, while the utilization of multiple scanning techniques helps establish a global receptive field, this approach does not effectively address the underlying causality issue. Tokens at different locations are still treated inconsistently due to their varying positions within the sequences, inevitably leading to a prioritization of more recently scanned tokens. Consequently, this strategy struggles to preserve spatial relationships and inadvertently introduces undesired bias for vision tasks as a result of the scanning order. In contrast, the self-attention mechanism in Transformers consistently considers all tokens to capture dependencies, showcasing its superiority to Mamba.
Secondly, the incorporation of multiple scanning strategies leads to redundancy and sub-optimal performance. To enhance Mamba’s perception at different locations, numerous approaches incorporate different scanning techniques including different scanning directions, axes, continuity, and sampling. While incorporating more scanning strategies may increase the information captured in specific locations, it introduces a significant amount of redundant information within these scanning sequences. This redundancy can impact model efficiency and hinder effective knowledge extraction for vision tasks.
Consequently, designing an appropriate approach for Mamba to consistently and elegantly consider tokens at different locations remains a critical challenge in visual data processing.
5.3. In-context Learning
Transformer models exhibit in-context learning (ICL) capabilities after large-scale pre-training, enabling them to learn new tasks with a few demonstrations without further explicit training or fine-tuning. In the dynamic landscape of deep learning, ICL methodologies have evolved to address increasingly complex tasks across NLP, CV, and multi-modal domains. These methodological advancements are pivotal for pushing the limits of existing deep learning frameworks.
Currently, Transformer models are the primary large models possessing ICL capabilities. Mamba’s efficiency and long sequence modeling ability make it a promising alternative to Transformers for ICL. Furthermore, the inherent causality of the Mamba model eliminates the need for position embeddings, which are typically a constraint in extending the ICL capabilities of Transformers. In (Park et al., 2024b), the authors investigated the ICL capabilities of Mamba compared to Transformers, focusing on regression ICL tasks and small-scale models. The results demonstrated that Mamba is capable of ICL and performs comparably to Transformers in standard regression ICL tasks. Additionally, Mamba outperforms Transformers in tasks such as sparse parity learning. However, Mamba struggles with tasks involving non-standard retrieval functionality, which Transformers can handle with ease. This limitation may stem from the memory-intensive nature of these tasks, as Mamba prioritizes efficiency over contextual expressivity by utilizing a shorter latent size (Jelassi et al., 2024). Mamba-2 demonstrates notable improvements over Mamba-1 in some of these tasks, though further investigation is needed to fully understand its performance (Dao and Gu, 2024a). Research detailed in (Grazzi et al., 2024) further demonstrated Mamba’s capabilities on more complex NLP in-context tasks and showed that Mamba’s ICL efficacy scales positively with the number of in-context examples.
With its proficient in-context modeling capabilities and adeptness at capturing long-range dependencies, the Mamba model shows promising potential for deeper semantic understanding and enhanced performance in ICL applications.
5.4. Trustworthiness
5.4.1. Interpretability
Some studies have provided experimental evidence to elucidate the mechanisms underlying the Mamba model in NLP, focusing on its in-context learning capabilities (Grazzi et al., 2024; Park et al., 2024b), factual recall capabilities (Sharma et al., 2024), and next-token prediction capabilities (Bachmann and Nagarajan, 2024). Additionally, other works (Cirone et al., 2024) have laid theoretical foundations for Mamba’s applications in NLP. Despite these advancements, explaining why Mamba performs effectively on visual tasks remains challenging (Jafari et al., 2024; Ali et al., 2024). In (Ali et al., 2024), the authors reinterpret Mamba layers as self-attention mechanisms, thereby shedding light on its functionality both empirically and theoretically. However, the distinct learning characteristics of visual Mamba and its parallels with other foundational architectures, such as RNNs, CNNs, and ViTs, still demand deeper interpretation.
5.4.2. Generalization
According to findings in (Long et al., 2024; Guo et al., 2024b), the hidden states in Mamba are likely to accumulate or even amplify domain-specific information, which can adversely affect its generalization performance. Moreover, the 1D scanning strategies inherent to the model may inadvertently capture domain-specific biases, and current scanning techniques often fail to address the need for domain-agnostic information processing. Addressing these challenges to better harness Mamba’s strengths while improving its generalization remains a critical area for future research.
5.4.3. Safety
Research presented in (Du et al., 2024a) demonstrated VMamba’s (Liu et al., 2024e) strengths in adversarial resilience and general robustness. However, it also identified limitations in scalability when dealing with these tasks. The study includes white-box attacks on VMamba to examine the behavior of its novel components under adversarial conditions. The findings indicated that while parameter $\Delta$ exhibits robustness, parameters $\mathbf{B}$ and $\mathbf{C}$ are susceptible to attacks. This differential vulnerability among parameters contributes to VMamba’s scalability challenges in maintaining robustness. Furthermore, the results revealed that VMamba exhibits particular sensitivity to interruptions in the continuity of its scanning trajectory and the integrity of spatial information. Enhancing the safety of visual Mamba remains an unresolved challenge.
6. Conclusion
Mamba has rapidly emerged as a transformative long sequence modeling architecture, renowned for its exceptional performance and efficient computational implementation. As it continues to gain significant attraction in the field of computer vision, this paper offers a comprehensive review of visual Mamba approaches. We begin with an in-depth overview of the Mamba architecture, progressing to detailed examinations of representative visual Mamba backbone networks and their extensive applications. These applications are systematically categorized by differing modalities, including image, video, point cloud, and multi-modal data. Finally, we analyze the challenges and delineate future directions for visual Mamba, providing valuable outlooks that may influence ongoing and future developments in this dynamically evolving field.
Acknowledgements.
This work was supported by the Hong Kong Innovation and Technology Fund (Project No. MHP/002/22), Project of Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone (HZQB-KCZYB-2020083) and the Research Grants Council of the Hong Kong (Project Reference Number: T45-401/22-N).